### v1

Abstract
Regular black holes are often used to address singularities, but they typically involve a potentially unstable de Sitter core and an internal horizon that breaks predictability. Integrable singularities (IS) have recently gained attention because they avoid both issues and exhibit finite tidal forces, allowing nondestructive radial infall. First, we present a new BH solution sourced by a string fluid (FS) that exhibits an IS. Motivated by the divergence of the conserved energy in the cloud of strings (CS) model, we introduce an energy density profile based on the screening of the CS energy density within an FS framework, yielding a finite conserved energy. On the other hand, it has been proposed Ovalle (2024) that an interior region, rather than a pointlike mass, can generate a Schwarzschild BH exterior region. Secondly, motivated by the variety of BH solutions with singularities in the literature, we establish the conditions that an interior region with an IS must satisfy to represent a generic exterior BH solution, with Schwarzschild being only a particular case of the latter. We derive the junction conditions (JC) between the interior and exterior regions, showing that they lead to temperature continuity at the interface, while discontinuities in tangential pressure lead to phase transitions. We propose that the nature of the interior region is described by CS and FS, while the exterior corresponds to Reissner–Nordström.
I Introduction
The observation of gravitational waves produced by the merger of two spinning black holes Abbott et al. (2016) has firmly positioned black holes among the most compelling phenomena in gravitational physics. In general, black hole solutions are characterized by the presence of a central singularity, where the spacetime description breaks down, signaled by the divergence of the curvature invariants. To address this problem, the standard approach involves specifying a particular form of matter within the energy-momentum tensor, which gives rise to regular black holes (RBHs). It is well established that RBH solutions typically include an inner horizon, which can also lead to the presence of a de Sitter–like core near the origin. In this context, several studies have indicated that the inner horizon is inherently unstable Poisson and Israel (1990); Brown et al. (2011). Along the same lines, a more recent study, reference Carballo-Rubio et al. (2023), argues that instabilities at the cores of regular black holes associated with the inner horizon are inevitable, since the mass inflation phenomenon represents a fundamental feature of RBHs with astrophysical significance. On the other hand, it is also worth noting that, as claimed in reference Ovalle (2023), predictability breaks down after crossing the inner horizon.
As an alternative to RBH solutions, Ref. Lukash and Strokov (2013) proposes a gravitational scenario in which a particle can reach the radial origin without being destroyed. Rather than a potentially unstable de Sitter core associated with an inner horizon, this model exhibits an integrable singularity at the origin. Integrable singularities are characterized by a divergence of the Ricci scalar while their volume integrals remain finite near the center. The elimination of the inner horizon thus prevents possible instabilities. Furthermore, recent studies in Ref. Estrada et al. (2024) show that a central integrable singularity allows for finite tidal forces, extendable radial geodesics, and a weak singularity, as formalized by Tipler’s extension theorem Nolan (2000). This ensures that an object approaching the central singularity would not undergo spaghettification. In addition, Ref. Casadio et al. (2023) offers a quantum perspective on the existence of such integrable singularities. For other recent applications, see Refs. Ovalle (2023); Estrada (2025).
On the other hand, String theory posits that the fundamental constituents of the universe are one-dimensional objects called strings Younesizadeh and Jokar (2022). The dynamics of a single 1-brane string can be extended to an ensemble Yang et al. (2023), forming a cloud of strings. Embedding such a cloud of strings modifies the spacetime geometry, producing black hole solutions surrounded by cloud of strings Letelier (1979), and several recent studies have analyzed black holes interacting with string clouds Zafar et al. (2025); Javed et al. (2026); Gogoi et al. (2025); Ahmed et al. (2025); Cai et al. (2025). A generalization of this model includes pressure or tension in the energy–momentum tensor, giving it a structure similar to a perfect fluid, i.e., a fluid of strings Letelier (1983). In this connection, recently an equation of state $\rho(r)=\alpha(r)p(r)$ was proposed Nunes dos Santos (2025) for a fluid of strings, leading to a regular black hole solution.
Following Ref. Aoki et al. (2021), where the volume integral $\int_{0}^{\infty}4\pi r^{2}T^{0}{}_{0},dr$ is interpreted as the conserved energy associated with a timelike Killing vector, we note that this integral diverges for the standard string-cloud energy density. Motivated by this, as the first problem in this work we introduce a modified version of the string-cloud energy density within the fluid-of-strings framework. This modification can be interpreted as a geometrical screening of the original distribution. As a result, we show that the volume integral converges to a finite value equal to $M$, which can be interpreted as the conserved energy charge. In this way, we provide a new fluid-of-strings black hole, constructed through the equation of state $\rho(r)=\alpha(r)p(r)$. Furthermore, we will show that the Ricci scalar near the origin behaves in a way consistent with an integrable singularity, and that the new solution is also free of a potentially unstable inner horizon.
On the other hand, the Schwarzschild solution, which according to Birkhoff’s theorem Hawking and Ellis (2023) is the unique spherically symmetric vacuum solution of Einstein’s equations, possesses a strong central singularity. In this paragraph, we point out some ideas discussed in Ref. Ovalle (2024). The Schwarzschild black hole features an event horizon, which can be viewed as a membrane of radius $h$ separating two causally disconnected space-time regions: an interior and an exterior one. The interior region, defined by $r<h$, contains the central singularity shielded by the event horizon. In this domain, the metric function $f(r)$ becomes negative, causing the radial and temporal coordinates to interchange their roles and thereby revealing the dynamical character of the interior geometry. As a consequence, gravitational collapse toward the central singularity becomes unavoidable, in agreement with Penrose’s theorem Penrose (1965). Conversely, an observer in the exterior region has no causal access to the interior of the black hole: once the boundary at $r=h$ is crossed, it is impossible to return or communicate with the exterior. This Ref. Ovalle (2024) examines whether an alternative structure, beyond a pointlike mass, can give rise to the Schwarzschild black hole region for $r\geq h$. In particular, the authors aim to describe an interior configuration associated with this region while keeping the exterior Schwarzschild geometry unchanged and ensuring that tidal forces remain finite everywhere. In connection with the latter, the authors propose that the internal region may contain either a central singularity or a regular core endowed with a potentially unstable inner horizon.
It is well known that a wide class of black hole solutions studied in the literature are supported by matter sources that inevitably lead to a central singularity, for example BH cosmic-void density profiles Lustosa et al. (2025), BH dark-matter with dark matter energy profile Xu et al. (2018), and hairy BHs Ovalle et al. (2021). Electrically charged BH, exemplified by the Reissner–Nordström solution, also arise from electromagnetic sources in the energy–momentum tensor and are characterized by the presence of an inner horizon. Motivated by the above considerations, and in analogy with Ref. Ovalle (2024), we address as a second problem the existence of an internal geometry inside the event horizon of a generic black hole spacetime, keeping the external geometry completely general, with the Schwarzschild solution appearing only as a particular case. We focus on scenarios in which the internal structure is endowed with an integrable singularity, in order to avoid the emergence of a potentially unstable de Sitter core and the presence of an inner horizon, since predictability breaks down after crossing the inner horizon. We investigate the possible nature of matter sources capable of supporting such an internal geometry with an integrable singularity, specifically those associated with clouds of strings and with the new string fluid introduced in this work. To do this, we define the conditions that the internal geometry must satisfy in order to represent an integrable singularity, as well as the junction conditions between generic internal and external black hole geometries. Furthermore, we examine the connection between the Israel–Darmois Israel (1966) junction conditions and black hole thermodynamic properties, and explore their relation to the existence of phase transitions. As a test case, we study the Reissner–Nordström solution as an internal geometry, considering clouds of strings and the new string fluid as the first and second examples, respectively.
II A new Fluid of strings black hole solution with an integrable singularity
In this section, we first present a new string fluid solution, showing that it can represent a black hole solution over the entire domain $r\in[0,\infty)$.Below, in line with this work, we characterize it as the interior region of a Reissner–Nordström black hole. Reference Letelier (1983) proposes a generalization of the cloud of strings model by incorporating pressure or tension into the energy–momentum tensor. This model is the so-called fluid of strings (fs). See Appendix B for further details. Due to the symmetries of the spacetime and in connection with the structure of the energy–momentum tensor (70), its general form is given by $T^{t}_{t}=T^{r}_{r}=-\rho^{(fs)}\,,\,T^{\theta}_{\theta}=T^{\phi}_{\phi}=p^{(fs)}$. It is of interest to establish an equation of state for the string fluid. In Ref. Nunes dos Santos (2025), an equation of state was proposed in which the energy density and the pressure are related through the form $\rho(r)^{(fs)}=\alpha(r)\cdot p(r)^{(fs)}$. In this reference, an equation of state with $\alpha(r)$ was proposed such that the solution of Einstein’s equations represents a regular string-fluid geometry at the origin. Within this framework, the energy–momentum tensor takes the form
$(T^{\mu}_{\ \nu})^{(fs)}=\mathrm{diag}\left(-\rho(r)^{(fs)},-\rho(r)^{(fs)},\dfrac{\rho(r)^{(fs)}}{\alpha(r)},\dfrac{\rho(r)^{(fs)}}{\alpha(r)}\right)$ (1)
We consider a line element
$ds^{2}=-f(r)_{(fs)}\,dt^{2}+\frac{dr^{2}}{f(r)_{(fs)}}+r^{2}d\Omega^{2}.$ (2)
As mentioned above, in this section the radial coordinate extends over the entire domain $r\in[0,\infty)$ and, consequently, there is an asymptotically flat boundary. As indicated in Ref. Nunes dos Santos (2025), within this framework the Einstein equations lead to a solution of the form:
$f(r)_{(fs)}=1+\frac{c_{2}}{r}+\frac{c_{1}}{r}\,I(r),$ (3)
$I(r)=\int dr\,\exp\left(\frac{-2}{r\alpha(r)}\right)$ (4)
[…]
$\rho^{(fs)}=\frac{M}{4\pi b^{2}r^{2}}\exp(-r/b)$ (5)
where $b>0$ is a constant. Thus, the factor $\exp(-r/b)$ can be regarded as a form of geometrical screening of the cloud of strings energy density, i.e., $\rho_{cs}\sim\text{Const}/r^{2}\to\rho^{(fs)}\sim(\text{Const}/r^{2})\exp(-r/b)$. We can verify that the spatial integral of the energy density is finite
$\int_{0}^{\infty}4\pi r^{2}\rho^{(fs)}\,dr=M\Rightarrow\mbox{finite}$ (6)
Thus, following the definition and assumptions of Aoki et al. (2021), our energy density, which screens the cloud-of-strings profile, yields a finite total energy, in contrast to the original cloud-of-strings profile.
The (t,t)(t,t)$(r,r)$ and tangential components of the equations of motion are:
$-\frac{\left(r(1-f_{(fs)})\right)^{\prime}}{r^{2}}=$ ${p}_{r}^{\,\,(fs)}=-{\rho}^{(fs)}$ (7)
$\frac{\left(r^{2}(f_{(fs)})^{\prime}\right)^{\prime}}{2r^{2}}=$ ${p}_{\theta}^{\,\,(fs)}$ (8)
Thus, by evaluating our energy density profile, the temporal and radial components of the Einstein equations lead to the following solution
$f(r)_{(fs)}=1-\frac{2M}{r}\left(1-\exp(-r/b)\right)$ (9)
By comparing Eqs. (3) and (9), it is straightforward to note that $c_{2}=-2M$ and, moreover, that:
$2M\,\exp\!\left(-\frac{r}{b}\right)=c_{1}\int dr\,\exp\!\left(-\frac{2}{r\,\alpha(r)}\right)$
- $\frac{2M}{b}\exp(-r/b)=c_{1}\,\exp\!\left(-\frac{2}{r\,\alpha(r)}\right)$
$\Rightarrow c_{1}=-\frac{2M}{b}\,\,\land\,\,\alpha(r)=\frac{2b}{r^{2}}$ (10)
[…]
$(T^{3}_{3})^{(fs)}=(T^{4}_{4})^{(fs)}=p_{\theta}^{\,\,(fs)}=\frac{M}{8\pi b^{3}}\exp(-r/b)$ (11)
II.1 Integrable singularity
The trace of the energy–momentum tensor (1) is given by $T^{(fs)}=-2\rho^{(fs)}+2p_{\theta}^{\,\,(fs)}=-R$, where $R$ is the Ricci scalar. Thus, we obtain the following relation:
$R=-T^{(fs)}=\frac{M}{2\pi b^{2}r^{2}}\exp(-r/b)-\frac{M}{4\pi b^{3}}\exp(-r/b)$ (12)
$\Rightarrow R\big|_{r\to 0}\sim r^{-2}$ (13)
Furthermore, replacing Eqs. (7) and (8), we obtain
$r^{2}\cdot R=2\left(r(1-f_{(fs)})\right)^{\prime}-{\left(r^{2}(f_{(fs)})^{\prime}\right)^{\prime}}$ (14)
We note that although the Ricci scalar diverges near the origin, see Eq. (13), the equations of motion (28) are integrable and remain free of singularities. This situation can therefore be described as an integrable singularity.
II.2 Absence of an inner horizon
First, we note that function (9) behaves near the origin as $f_{(fs)}|_{r\sim 0}\sim 1-2M/b$, with $2M/b>1$ in order for the metric signature to be +,,,+,-,-,- near the origin. Moreover, $|1-\frac{2M}{b}|$ is sufficiently large and dominant in the vicinity of the origin, so as to avoid the appearance of a zero of the function $f(r)_{(fs)}$ near the origin. We observe the numerical behavior and the absence of a potentially unstable inner horizon in Fig. 1 . Consequently, our solution lacks the presence of an unstable de Sitter core, unlike other black hole solutions sourced by matter.
Figure 1: $f(r)_{(fs)}$ for $b=1,M=0.8,0.9,1$. We observe the absence of a potentially unstable inner horizon.
On the other hand, we note that the derivative $df_{(fs)}/dr=2M/r^{2}\left(1-\exp(-r/b)(1+r/b)\right)$ is always positive. Consequently, since $f(0)_{(fs)}=f_{\min}<0$, there is only a single sign change, from negative to positive, at the event horizon. For both reasons discussed above, we can conclude that there is no presence of a potentially unstable inner horizon.
In this way, in this section we have presented a new string fluid solution that represents a geometry with an integrable singularity and the absence of a potentially unstable inner horizon. As mentioned above, in this section we analyze a black hole solution over the entire domain $r\in[0,\infty)$. Below, in line with this work, we characterize it as the interior region of a Reissner–Nordström black hole.
III Black Hole Solutions with an Interior Geometry Featuring an Integrable Singularity
In Fig. 2, we schematically display the configuration of our geometry, which consists of two regions: an interior region, where the radial coordinate runs from the origin to the event horizon at $r=h$, i.e. $r\in[0,h]$, and an exterior region, extending from the event horizon to infinity (or to a cosmological horizon, depending on the case). Below, we describe both regions.
hr
Figure 2: Schematic representation of the interior region $r\in[0,h]$ and the exterior region extending from $r=h$ to the black hole boundary.
III.1 Description of the interior spacetime :
It corresponds to a static and spherically symmetric spacetime, where the radial coordinate runs over $r\in[0,h]$. The line element is given by:
$ds^{2}=-f(r)\,dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}d\Omega^{2}.$ (15)
where $h$ corresponds to the event horizon, such that $f(h)=0$. Since this region represents the interior of the black hole, the signature of the metric tensor (15) is +,,,+,-,-,-, and therefore $f(r)\leq 0$ in this region. In this way, $f(r)$ is an increasing function, from its minimum value $f(r=0)=f_{\min}<0$ up to $f(h)=0$. First, we consider the presence of a cosmological constant in this region, where the matter sources in the interior are described by the interior energy–momentum tensor.
$(T^{\mu}_{\ \nu})^{(in)}=\mathrm{diag}\!\left(-\rho,\,p_{r},\,p_{\theta},\,p_{\theta}\right)$ (16)
The (t,t)(t,t)$(r,r)$ and tangential components of the equations of motion are:
$-\frac{\left(r(1-f)\right)^{\prime}}{r^{2}}+\Lambda=$ $-\rho=p_{r}$ (17)
$\frac{\left(r^{2}f^{\prime}\right)^{\prime}}{2r^{2}}+\Lambda=$ $p_{\theta}$ (18)
where we observe that $\rho=-p_{r}$. The conservation equation reads:
Thus, we can define the effective interior energy-momentum tensor $(\bar{T}^{\mu}_{\ \nu})^{(\text{in})}$, for $\Lambda=-3/l^{2}$, where $l$ denotes the AdS radius.
$(\bar{T}^{\mu}_{\ \nu})^{(\text{in})}=\mathrm{diag}\!\left(-\bar{\rho},\,\bar{p}_{r},\,\bar{p}_{\theta},\,\bar{p}_{\theta}\right)$ (19)
Thus, the effective density and pressures are given by
$\bar{\rho}=$ $\rho+\Lambda=\rho-\frac{3}{l^{2}}$ (20)
$\bar{p}_{r}=$ $p_{r}-\Lambda=\frac{3}{l^{2}}+p_{r}=\frac{3}{l^{2}}-\rho=-\bar{\rho}$ (21)
$\bar{p}_{\theta}=$ $p_{\theta}-\Lambda=\frac{3}{l^{2}}+p_{\theta}$ (22)
In this way, the equations of motion can be written as:
$-\frac{\left(r(1-f)\right)^{\prime}}{r^{2}}=$ $\bar{p}_{r}=-\bar{\rho}$ (23)
$\frac{\left(r^{2}f^{\prime}\right)^{\prime}}{2r^{2}}=$ $\bar{p}_{\theta}$ (24)
III.2 Description of the exterior spacetime
It corresponds to a static and spherically symmetric spacetime, where the radial coordinate runs over $r\in[h,r_{b}]$, with $r_{b}$ denoting the boundary of the spacetime. In the presence of a cosmological horizon, $r_{b}\to r_{c}$, where $r_{c}$ is the cosmological horizon. In the absence of the latter, i.e., in asymptotically flat or asymptotically AdS spacetimes, $r_{b}\to\infty$. The line element is given by:
$ds^{2}=-f(r)_{E}\,dt^{2}+\frac{dr^{2}}{f(r)_{E}}+r^{2}d\Omega^{2}.$ (25)
The geometry is such that $f_{E}(h)=0$. Since this region represents the exterior of the black hole, the signature of the metric tensor (25) is ,+,+,+-,+,+,+, and therefore $f_{E}(r)\geq 0$ in this region. We consider the case without a cosmological constant, with an energy–momentum tensor of the form
$(T^{\mu}_{\ \nu})^{(E)}=\mathrm{diag}\left(-\rho^{(E)},p_{r}^{(E)},p_{\theta}^{(E)},p_{\theta}^{(E)}\right)$ (26)
It is straightforward to observe that the equations of motion can be obtained by replacing $f\to f_{E}$, $\bar{\rho}\to\rho^{(E)}$, $\bar{p}_{r}\to p_{r}^{(E)}$, and $\bar{p}_{\theta}\to p_{\theta}^{(E)}$ in Eqs. (23) and (24).
IV INTERIOR REGION AS AN INTEGRABLE SINGULARITY
In this section, we state the properties that a generic interior solution must satisfy in order to represent an integrable singularity near the origin. In the following sections, we analyze the specific cases in which the interior regions correspond, respectively, to a string cloud and to the new string-fluid solution provided in Section II.
The trace of our efective interior energy–momentum tensor (19) is given by $\bar{T}^{(in)}=-2\bar{\rho}+2\bar{p}_{\theta}$. Replacing Eqs. (23) and (24), we obtain
$R=\bar{T}^{(in)}=-\frac{2\left(r(1-f)\right)^{\prime}}{r^{2}}+\frac{\left(r^{2}f^{\prime}\right)^{\prime}}{2r^{2}}$ (27)
where $R$ is the Ricci scalar. Analogously to Section II, this equation corresponds to the trace of the Einstein equations and can be rewritten as
$r^{2}\cdot R=2\left(r(1-f)\right)^{\prime}-{\left(r^{2}f^{\prime}\right)^{\prime}}$ (28)
Some remarks:
  1. 1.
    We are interested in a function $f(r)$ that is finite at the origin and has no zeros in the interval $r\in[0,h]$. In this way, near the origin:
    $f(r\sim 0)\sim 1-a+\mathcal{O}(r^{n})\Rightarrow R\sim\frac{\mbox{Cte}}{r^{2}}$ (29)
    where $a>1$ and |1a||1-a| is sufficiently large and dominant compared to the term $\mathcal{O}(r^{n})$ near the origin, in order to avoid a zero of the function $f(r)$ in the vicinity of the origin. As mentioned above, the absence of zeros in this interval implies the absence of a potentially unstable inner horizon. Although the Ricci scalar has a singularity $\sim r^{-2}$ at the origin, the equations of motion (28) are integrable, being free of singularities, and lead to a finite solution $f(r)$ at the origin. This situation can be referred to as an integrable singularity. Under these assumptions, one can also observe that the Kretschmann scalar exhibits a singularity at $r=0$.
    $K=f^{\prime\prime}(r)^{2}+\frac{4f^{\prime}(r)^{2}}{r^{2}}+\frac{4(f(r)-1)^{2}}{r^{4}}$ (30)
  2. 2.
    As mentioned, $f(r)$ is an increasing function from its minimum value $f(r=0)=f_{\min}<0$ up to $f(h)=0$. In this way, $df/dr>0$ for $r\in[0,h]$, with the temperature $T=\left.df/dr\right|_{h^{-}}=\left.df_{E}/dr\right|_{h^{+}}$.
V Junction conditions
To construct a model describing a physically viable compact object, it is necessary to impose a smooth matching between the interior geometry and the exterior spacetime. For this purpose, we will consider the Israel–Darmois matching conditions Israel (1966).
  1. 1.
    At the surface $\Sigma$, identified with the event horizon radius $r=h$, these conditions imply the continuity of the geometry, known as the first fundamental form
    $ds^{2}]_{\Sigma(r=h)}=0\Rightarrow f(h)=f_{E}(h)=0$ (31)
    where $F]_{\Sigma\,\equiv\,r=h}\equiv F(h^{+})-F(h^{-})$.
  2. 2.
    The second fundamental form is given by $G_{\mu\nu}x^{\nu}]_{\Sigma(r=h)}=0$, where $x^{\nu}$ is a unit vector projected along the radial direction. This implies the following
    $\bar{p}_{r}(h)=$ $p_{r}^{(E)}$
    $\frac{1-hf^{\prime}(h)}{h^{2}}=$ $\frac{1-h\cdot\big(f_{E}(h)\big)^{\prime}}{h^{2}}$
    $\Rightarrow$ $f^{\prime}(h)=\big(f_{E}(h)\big)^{\prime}$
    $\Rightarrow$ $T_{in}=T_{E}=T$ (32)
    Since the derivative of $f^{\prime}(h)$ matches the derivative of the exterior region at the event horizon $\big(f_{E}(h)\big)^{\prime}=T$, in this work we also refer to the derivative of $f^{\prime}(h)$ as the temperature. Thus, the second fundamental form of the Israel–Darmois junction conditions leads to the continuity of the temperature between the interior and exterior regions.
  3. 3.
    As we will see below, this condition is not mandatory, unlike the two previous ones. More specifically, this condition represents a consequence of the discontinuity (or continuity) of the tangential pressure in black hole thermodynamics.We obtain the interior temperature from Eq. (24). In an analogous way, we obtain the exterior temperature for the corresponding region. Thus, we evaluate Eq. (2)
    $T_{in}=$ $T_{E}$
    $\cancel{h}\left(\bar{p}_{\theta}(h)-\frac{f^{\prime\prime}(h)}{2}\right)=$ $\cancel{h}\left(p_{\theta}^{(E)}(h)-\frac{\big(f^{\prime\prime}(h)\big)_{E}}{2}\right)$
    $\frac{3}{l^{2}}+p_{\theta}(h)-\frac{f^{\prime\prime}(h)}{2}=$ $p_{\theta}^{(E)}(h)-\frac{\big(f^{\prime\prime}(h)\big)_{E}}{2}$ (33)
    We use
    $C_{in}=T_{in}\frac{dS}{dh}\left(\frac{dT}{dh}\right)^{-1}=T_{in}\frac{dS}{dh}\left(\frac{d}{dh}\,\left(\frac{df}{dh}\right)\right)^{-1}$ $\Rightarrow f^{\prime\prime}(h)=\frac{T}{C_{in}}\frac{dS}{dh}$ (34)
    analogously $\Rightarrow\big(f^{\prime\prime}(h)\big)_{E}=\frac{T}{C_{E}}\frac{dS}{dh}$ (35)
    We assume that the entropy follows the area law, i.e. $S=\pi h^{2}$, thus evaluating (33)
    $\frac{C_{\text{in}}-C_{E}}{C_{\text{in}}\cdot C_{E}}=$ $\frac{1}{T\pi h}\,\left(p_{\theta}^{(E)}(h)-\left(\frac{3}{l^{2}}+p_{\theta}(h)\right)\right)$
    $\Rightarrow C_{\text{in}}-C_{E}\sim$ $\,\,\,p_{\theta}^{(E)}(h)-\bar{p}_{\theta}(h)$ (36)
    thus
    $\mbox{if}\,\,\bar{p}_{\theta}(h)=p_{\theta}^{(E)}(h)\Rightarrow$ $C_{\text{in}}=C_{E}$ (37)
    $\mbox{if}\,\,\bar{p}_{\theta}(h)\neq p_{\theta}^{(E)}(h)\Rightarrow$ $C_{\text{in}}\neq C_{E}$ (38)
    In our case the entropy is continuous. In this way, we test that the first derivative of the Gibbs potential, S=dG/dTS=-dG/dT, is also continuous. On the other hand, using $C=T\,dS/dT=-\,T\,d^{2}G/dT^{2}$, we observe that a discontinuity in the tangential pressure implies the existence of a second-order phase transition.
VI Example 1: A Reissner–Nordström Exterior with a Cloud of Strings as the Interior Region with an Integrable Singularity
In Section  III, we described our geometric setup to represent both the interior and exterior regions of a generic black hole geometry. In Section IV, we formulated the conditions that an interior region must satisfy in order to represent an integrable singularity near the origin, while in Section V we presented the junction conditions between the interior and exterior regions. However, the criteria introduced above must be tested in explicit black hole geometries. In this spirit, we analyze two illustrative examples. In the present section, we study a Reissner–Nordström exterior with a cloud of strings as the interior region. In the following section, we consider a Reissner–Nordström exterior with the new string-fluid solution provided in Section  II.
Remarks on the Reissner–Nordström (RN) Black Hole:
the electromagnetic field has an energy-momentum (EM) tensor that is symmetric and traceless, constructed from the Maxwell field tensor $F_{\mu\nu}=\nabla_{\mu}A_{\nu}-\nabla_{\nu}A_{\mu}$. The Maxwell equations are given by $\nabla_{\mu}F^{\mu\nu}=0$ and $\nabla_{[\gamma}F_{\mu\nu]}=0$. Solving the Einstein–Maxwell equations for a static, spherically symmetric electric charge leads to the RN metric. In this case, the metric function $f_{E}(r)$ appearing in the line element (25), expressed in natural units, is given by
$f(r)_{E}=1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}$ (39)
where $M,Q$ correspond to the mass and electric charge, respectively. The inner and event horizons, $r_{-}$ and $h$, are given, respectively, by:
$r_{-}=M-\sqrt{M^{2}-Q^{2}}$ (40)
$h=M+\sqrt{M^{2}-Q^{2}}$ (41)
Thus, we note that it must be satisfied that $M>Q\Rightarrow M^{2}-Q^{2}>0$. As we have mentioned, in our setup, the exterior geometry extends over $r\in[h,r_{b}]$, with the boundary corresponding to the radial coordinate approaching infinity, since the solution is asymptotically flat. Therefore, we are interested in the aforementioned range of the radial coordinate. The temperature is given by:
$T=\frac{1}{4\pi}\left.\frac{df_{E}}{dr}\right|_{r=h}=\frac{1}{4\pi}\frac{2\sqrt{M^{2}-Q^{2}}}{\left(M+\sqrt{M^{2}-Q^{2}}\right)^{2}}$ (42)
The exterior energy–momentum tensor (26) takes the form:
$(T^{\mu}_{\ \nu})^{(E)}=\mathrm{diag}\left(-\rho^{(E)}=-E^{2}(r),p_{r}^{(E)}=-E^{2}(r),p_{\theta}^{(E)}=E^{2}(r),p_{\theta}^{(E)}=E^{2}(r)\right)$ (43)
For our line element, the electric field is given by
$E(r)=\sqrt{-F_{01}F^{01}}=\frac{Q}{r^{2}}\,.$ (44)
VI.1 Representation of a Cloud of Strings as an Interior Region with an Integrable Singularity
In this subsection, we explore a possible nature of the matter sources that could give rise to an interior region characterized by an integrable singularity. Specifically, we consider a cloud of strings as a potential matter source. In this context, as previously mentioned, the embedding of an ensemble of 1-branes—the so-called cloud of strings—can deform the spacetime geometry Letelier (1979). See Appendix A for a more detailed description of a cloud of strings.
The interior energy–momentum tensor $(T^{\mu}_{\ \nu})^{(in)}$, equations (16), reported in reference Letelier (1979), Eq. (68), has the form
$(T^{\mu}_{\ \nu})^{(in)}=\mbox{diag}\left(-\rho,-\rho,0,0\right)$ (45)
where the energy density originates from Eq. (69) has the form
$\rho=\frac{a}{r^{2}}$ (46)
Thus, from equations (20),(21), (22), we note that the interior energy density and effective pressures take the following form
$\bar{\rho}=-\bar{p}_{r}=\frac{a}{r^{2}}-\frac{3}{l^{2}}\,\,,\,\,\bar{p}_{\theta}=\frac{3}{l^{2}}$ (47)
Thus, we can observe that the trace of the effective interior energy-momentum tensor $(\bar{T}^{\mu}_{\ \nu})^{(\text{in})}$, given by equations (19), (20), (21) and (22) is
$\bar{T}^{(\text{in})}=-2\bar{\rho}+2\bar{p}_{\theta}=-\frac{2a}{r^{2}}+\frac{12}{l^{2}}=-R.$ (48)
Thus, we note that the energy-momentum tensor behaves as $R\sim r^{-2}$ at short length scales. This implies that the equations of motion (28) are integrable near the origin, being free of singularities. Thus, condition IV 1 is satisfied for the string cloud model to represent the interior region as an integrable singularity.
Below, we present a solution to the equations of motion characterized by the sources of a cloud of strings together with a cosmological constant, as given in Eq. (47). The solution of the equations of motion (23) and (24), together with the trace equation (28), gives rise to the following metric function $f(r)$, which remains finite at the origin
$f(r)=1-a+\frac{r^{2}}{l^{2}}$ (49)
where $a>1$ and |1a||1-a| is sufficiently large compared to the term $r^{2}/l^{2}$ near the origin, in order to avoid a zero of the function $f(r)$ in the vicinity of the origin. We have set the integration constants to zero since they lead to a singularity in the metric tensor. This result is further reinforced by the fact that, in this work, we are interested in an interior geometry associated with an integrable singularity, for which the Ricci tensor is such that the equations of motion remain integrable.
In order to test condition IV 2 for the interior region to be represented as an integrable singularity, we note that $f(r)$ is an increasing function, starting from its minimum value f(r=0)=1a<0f(r=0)=1-a<0, with $a>1$, and reaching $f(h)=0$. Indeed, $df/dr=2r/l^{2}>0$ for $r\in[0,h]$, and therefore the function is monotonically increasing and negative throughout this interval. Therefore, we point out that the parameter $a$ must satisfy $a>1$ in order for this condition to be fulfilled.
VI.2 Junction conditions:
As mentioned in Equation (41), the event horizon of the external RN region is given by $h=M+\sqrt{M^{2}-Q^{2}}$.
  • In order to satisfy junction condition V.1, i.e. the so-called first fundamental form of the Israel–Darmois formalism, upon evaluating the functions (39) and (49) on (31), one obtains the following condition
    $a=1+\left(\frac{M+\sqrt{M^{2}-Q^{2}}}{l}\right)^{2}$ (50)
    which is consistent with the previously mentioned condition that $a>1$.
  • In order to satisfy the junction condition V.2, i.e., the so-called second fundamental form of Israel–Darmois, and by using Equation (2), the continuity of the temperature at the event horizon $r=h$ leads to the following value of the cosmological constant in the interior region:
    $\Lambda=-\frac{3}{l^{2}}=-\frac{3\sqrt{M^{2}-Q^{2}}}{\left(M+\sqrt{M^{2}-Q^{2}}\right)^{3}}$ (51)
  • As mentioned, the third condition III.3 is not mandatory like the previous two, since it does not stem from the Israel–Darmois junction conditions. However, it is useful to test whether a phase transition occurs at the event horizon $r=h$. Using Equation (3):
    $C_{\text{in}}-C_{E}\sim\bar{p}_{\theta}(h)-p_{\theta}^{(E)}(h)=-\Lambda-\frac{Q^{2}}{h^{4}}=\frac{3M\left(M+\sqrt{M^{2}-Q^{2}}\right)-4Q^{2}}{\left(M+\sqrt{M^{2}-Q^{2}}\right)^{4}}.$ (52)
    Testing the denominator of the last equation for zero leads us to:
    $Q^{2}\left(\frac{16Q^{2}}{9M^{2}}-\frac{5}{3}\right)=0\Rightarrow Q_{c}^{2}=\frac{15}{16}M^{2}$ (53)
    where the previously mentioned condition $M^{2}>Q^{2}$ is satisfied for $Q=Q_{c}$. We have discarded $Q=0$, since this would lead to an external solution without electric charge. That is, for all values of $Q^{2}\neq Q_{c}^{2}$, a second-order phase transition occurs at $r=h$ for the parameters mentioned above. Conversely, for $Q^{2}=Q_{c}^{2}$, no phase transition occurs at this location.
VII Example 2: A Reissner–Nordström Exterior with a new fluid of Strings as the Interior Region with an Integrable Singularity
In Section II, a new string fluid solution was introduced. There, it was presented under the assumption that it extends over the entire spacetime, $r\in[0,\infty)$. In the present section, we analyze instead the case in which the solution of Section II, in the presence of a negative cosmological constant, is defined only in the interval $r\in[0,h]$, i.e., it can represent the interior region of the Reissner–Nordström solution, as depicted in Scheme 2. It is worth mentioning that, in Section VI, we have already highlighted the main features of the Reissner–Nordström solution that are relevant to this work.
VII.1 Representation of the New Fluid of Strings solution as an Interior Region with an Integrable Singularity
The effective interior energy–momentum tensor (19), obtained from Eqs. (20), (21), and (22), (5) and (11), takes the following form
$\bar{\rho}=-\bar{p}_{r}=\frac{M}{4\pi b^{2}r^{2}}\exp(-r/b)-\frac{3}{l^{2}}$ (54)
$\bar{p}_{\theta}=\frac{M}{8\pi b^{3}}\exp(-r/b)+\frac{3}{l^{2}}$ (55)
We can observe that the trace of the effective interior energy–momentum tensor $(\bar{T}^{\mu}_{\ \nu})^{(\text{in})}$, which includes the contribution of the cosmological constant, is given by
$\bar{T}^{(\text{in})}=-2\bar{\rho}+2\bar{p}_{\theta}=-\frac{M}{2\pi b^{2}r^{2}}\exp(-r/b)+\frac{M}{4\pi b^{3}}\exp(-r/b)+\frac{12}{l^{2}}=-R.$ (56)
Thus, the trace of the effective interior energy–momentum tensor with a cosmological constant behaves as $R\sim r^{-2}$ at short length scales. This behavior implies that the equations of motion (28) are integrable in the vicinity of the origin and remain free of pathological singularities. Consequently, condition IV 1 is satisfied, allowing this new string fluid model to consistently represent the interior region as an integrable singularity.
The solution is given by:
$f(r)=1-\frac{2M}{r}\left(1-\exp(-r/b)\right)+\frac{r^{2}}{l^{2}}$ (57)
which behaves as $f\mid_{r\sim 0}\sim 1-\frac{2M}{b}+\frac{r^{2}}{l^{2}}$ near the origin, with $\frac{2M}{b}>1$ and $|1-\frac{2M}{b}|$ sufficiently large compared to the term $\frac{r^{2}}{l^{2}}$ in the vicinity of the origin, so as to avoid a zero of the function $f(r)$ near the origin.
In order to test condition IV 2 for the interior region to be represented as an integrable singularity, we note that $f(r)$ is an increasing function, starting from its minimum value $f(r=0)=1-\frac{2M}{b}<0$, with $\frac{2M}{b}>1$, and, as we shall see below, reaching $f(h)=0$. Indeed, $df/dr=\frac{2M}{r^{2}}\left(1-\exp(-r/b)\left(1+\frac{r}{b}\right)\right)+\frac{2r}{l^{2}}>0$ for $r\in[0,h]$.
VII.2 Junction conditions:
In order to satisfy junction condition V 1, i.e. the so-called first fundamental form of the Israel–Darmois formalism, upon evaluating the functions (39) and (57) on (31), one obtains the following conditions:
$\frac{2M}{h}\exp(-h/b)=\frac{Q^{2}l^{2}-h^{4}}{h^{2}l^{2}}$ (58)
where the event horizon $h$ is given by Eq. (41). The previous equation, in turn, leads to the following:
$l^{2}>\frac{h^{4}}{Q^{2}}$ (59)
In order to satisfy the junction condition V 2, i.e., the so-called second fundamental form of Israel–Darmois, and by using Equation (2), the continuity of the temperature at the event horizon $r=h$ leads to the following condition:
$\frac{2M}{h}\exp(-h/b)=\frac{2b(Q^{2}l^{2}+h^{4})}{h^{2}l^{2}(b+h)}$ (60)
We can observe that the left-hand sides of Eqs. (58) and (60) coincide. By equating both equations, we obtain the following value for the AdS radius
$l^{2}=\frac{h^{4}}{Q^{2}}\cdot\left(\frac{3b+h}{h-b}\right)$ (61)
Thus, for this value of $l^{2}$, with $b,h>0$ we must evaluate condition (59)
$\frac{3b+h}{h-b}>1$ $\Rightarrow\frac{4b}{h-b}>0$
$\Rightarrow h>b$ (62)
where $h$ is given by Eq. (41). Thus, condition (62) must be satisfied.
As mentioned, the third condition V 3 is not mandatory like the previous two, since it does not stem from the Israel–Darmois junction conditions. However, it is useful to test whether a phase transition occurs at the event horizon $r=h$. Using Equation (3):
$C_{\text{in}}-C_{E}\sim\bar{p}_{\theta}(h)-p_{\theta}^{(E)}(h)=\frac{h}{16\pi b^{3}}\frac{2M}{h}\exp(-r/b)+\frac{3}{l^{2}}-\frac{Q^{2}}{h^{4}}$ (63)
using equations (58) and (61)
$C_{\text{in}}-C_{E}\sim\bar{p}_{\theta}(h)-p_{\theta}^{(E)}(h)=-\frac{Q^{2}(24\pi b^{3}-8\pi b^{2}h-h^{3})}{4\pi b^{2}h^{4}(3b+h)}$ (64)
Numerically, we find that the last equation vanishes for
$b=b_{c}=0.4116145346\cdot h$ (65)
Hence, there is no phase transition at $b=b_{c}$, whereas a phase transition occurs for $b\neq b_{c}$.
VIII Discussion and Summary
In Section II, we present a new fluid-of-strings (fs) black hole solution defined over the entire domain $r\in[0,\infty)$ and characterized by the equation of state $\rho^{(fs)}(r)$ and $p^{(fs)}(r)$. Following Ref. Aoki et al. (2021), where the integral $\int_{0}^{\infty}4\pi r^{2}T^{0}{}_{0},dr$ is interpreted as the conserved energy associated with a timelike Killing vector, we note that this integral diverges for the standard string-cloud energy density. Motivated by this behavior, we introduce a modified energy density for the fluid of strings, which can be interpreted as a geometrical screening of the original string-cloud distribution. As a result, the aforementioned integral converges to a finite value equal to $M$, which can be interpreted as the conserved energy charge. Although the Ricci scalar diverges near the origin as $R\sim r^{-2}$, these divergences are integrable, and the spacetime remains free of curvature singularities, allowing the solution to be classified as an integrable singularity. Finally, we show that the solution is free of a potentially unstable de Sitter core and of an inner horizon, whose presence is associated with a breakdown of causality.
Motivated by the fact that several black hole solutions sourced by matter fields leading to a destructive central singularity have attracted considerable interest in the literature, and in analogy with Ref. Ovalle (2024), we have addressed, as a second problem, the existence of an internal geometry inside the event horizon of a generic black hole spacetime. Within this framework, the external geometry is kept completely general, with the Schwarzschild solution appearing only as a particular case. We have focused on scenarios in which the internal structure is endowed with an integrable singularity, in order to avoid the emergence of a potentially unstable de Sitter core and the presence of an inner horizon, since predictability breaks down after crossing the inner horizon. We have investigated the nature of the matter sources capable of supporting such an internal geometry with an integrable singularity, specifically those associated with clouds of strings and with the new string fluid introduced in this work. To this end, in Section IV we have defined the conditions that the internal geometry must satisfy in order to represent an integrable singularity. In Section V we have established the junction conditions between the internal and generic external black hole geometries. We have shown that the second fundamental form of the Israel–Darmois junction conditions Israel (1966) leads to the continuity of the temperature across the interface between the internal and external regions. Furthermore, from the analysis of the junction conditions, we have determined that a discontinuity in the tangential pressure implies the existence of a phase transition at the location of the event horizon, and vice versa.
In Sections VI and VI, we studied the cases in which the internal geometry of the Reissner–Nordström spacetime is described, respectively, by a string cloud and by our newly proposed string fluid geometry. In both scenarios, the parameters of the corresponding internal geometries satisfy the conditions listed in Section IV, ensuring the presence of an integrable singularity near the origin, the absence of a potentially unstable de Sitter core, and the absence of an internal horizon that could lead to a breakdown of causality. In Section VI, we determined the allowed values of the energy density parameter $a$ and of the internal cosmological constant required to satisfy both Israel junction conditions at the interface between the external RN region and the internal string cloud region. In addition, we identified a critical value of the electric charge for which no phase transition occurs in the location of the interface between the two regions. In Section VI, we derived the conditions that the AdS radius, the electric charge, the screening parameter $b$, and the event horizon must satisfy in order to fulfill both Israel junction conditions at the interface between the external RN region and the internal region described by the new string fluid. Moreover, we numerically determined the critical value of the screening parameter for which no phase transition occurs in the location of the interface between the two regions.
Appendix A A brief revision of the cloud of string solution
In this appendix, we review the necessary tools for our study of the so-called cloud of strings from reference Letelier (1979). A 1-brane string is parametrized by the parameters $\lambda^{a}=(\lambda^{0},\lambda^{1})$, which is embedded into spacetime as $x^{\mu}=x^{\mu}(\lambda)$. The trajectory of a 1-brane defines a two-dimensional worldsheet. The Nambu–Goto (NG) action for a 1-brane is proportional to the area of this worldsheet.
$S_{NG}=\int\mathcal{M}\,dA=\int\mathcal{M}\sqrt{-h}d\lambda^{0}d\lambda^{0}\,\,;\,\,h_{ab}=g_{\mu\nu}\frac{\partial x^{\mu}}{\partial\lambda^{a}}\frac{\partial x^{\nu}}{\partial\lambda^{b}}$ (66)
Let $\mathcal{M}$ be a constant such that the action is dimensionless in natural units. Usually, $\mathcal{M}]=[T_{0}/c]=[\text{Force}]=[\text{Energy}/L]=[L^{-2}]$, where $T_{0}$ and $c$ are the 1-brane tension and the speed of light, respectively. The induced metric is given by $h_{ab}$, while $\sqrt{-h}$ corresponds to its determinant. This reference introduces a bivector $\Sigma^{\mu\nu}$ that spans the two-dimensional time-like worldsheet of the string. It can be expressed as
$\Sigma^{\mu\nu}=\epsilon^{AB}\frac{\partial x^{\mu}}{\partial\xi^{A}}\frac{\partial x^{\nu}}{\partial\xi^{B}},$ (67)
where $\epsilon^{AB}$ is the two-dimensional Levi-Civita symbol with components $\epsilon^{01}=-\epsilon^{10}=-1$. The action in (66) leads to an energy-momentum tensor for a 1-brane string. This reference shows that for an ensemble or cloud of strings, this tensor takes the form:
$T^{\mu\nu}=-\frac{\rho_{cs}\,\Sigma^{\mu\alpha}\Sigma_{\alpha}^{\;\;\nu}}{\sqrt{-h}}\,\,;\,\,h=\frac{1}{2}\Sigma_{\mu\nu}\Sigma^{\mu\nu}$ (68)
For the energy–momentum tensor reported in this reference, Eq. (68), the only nonzero components are $T^{0}_{0}=T^{1}_{1}$. The conservation equation $\nabla_{\mu}T^{\mu\nu}=0$ implies that the structure of this tensor is given by
$T^{0}_{0}=-T^{1}_{1}=\rho=\sqrt{-h}\,\rho_{cs}=\frac{a}{r^{2}},$ (69)
where $\rho_{cs}$ can be interpreted as the proper rest energy density of the string cloud.
Appendix B Some general aspects of the fluid of strings solution
As previously discussed, the original cloud of strings framework was subsequently extended in Ref. Letelier (1983) in order to incorporate the presence of an effective pressure. Within this generalized description, the energy–momentum tensor takes the form
$T^{\mu\nu}=\left(p+\rho_{cs}\sqrt{-h}\right)\frac{\Sigma^{\mu\lambda}\Sigma^{\nu}{}_{\lambda}}{(-h)}+p\,g^{\mu\nu},$ (70)
where $p$ and $\rho_{cs}$ denote, respectively, the pressure and the energy density associated with the string fluid. Considering a line element of the form (15). As indicated in Ref. Soleng (1995), due to the symmetries of the metric, the only nonvanishing components of $\Sigma^{\mu\lambda}$ are $\Sigma^{tr}$ and $\Sigma^{\theta\phi}$; therefore, the determinant of the induced metric satisfies $h<0$.

### v2

@@
 First, we present a new BH solution sourced by a string fluid (FS) that exhibits an IS.
-Motivated by the divergence of the conserved energy in the cloud of strings (CS) model, we introduce an energy density profile based on the screening of the CS energy density within an FS framework, yielding a finite conserved energy.
-On the other hand, it has been proposed Ovalle (2024) that an interior region, rather than a pointlike mass, can generate a Schwarzschild BH exterior region.
-Secondly, motivated by the variety of BH solutions with singularities in the literature, we establish the conditions that an interior region with an IS must satisfy to represent a generic exterior BH solution, with Schwarzschild being only a particular case of the latter.
+Motivated by the divergence of the conserved energy in the cloud of strings (CS) model, we introduce an energy density profile based on the screening of the CS energy density within an FS framework, thereby yielding finite conserved energy.
+On the other hand, the idea that an interior region, rather than a pointlike mass, can generate a Schwarzschild black hole exterior region has recently gained attention Ovalle (2024); Maeda (2025).
+This is obtained by matching an interior solution to the Schwarzschild exterior at the event horizon.
+Secondly, motivated by the variety of singular BH solutions in the literature, we establish the conditions that an interior region with an IS must satisfy in order to represent the interior of a generic exterior BH solution, with Schwarzschild being only a particular case of the latter.
 We derive the junction conditions (JC) between the interior and exterior regions, showing that they lead to temperature continuity at the interface, while discontinuities in tangential pressure lead to phase transitions.
@@
 For other recent applications, see Refs.
-Ovalle (2023); Estrada (2025).
+Ovalle (2023); Estrada (2025); Arrechea et al.
+(2025).
 On the other hand, String theory posits that the fundamental constituents of the universe are one-dimensional objects called strings Younesizadeh and Jokar (2022).
@@
 Conversely, an observer in the exterior region has no causal access to the interior of the black hole: once the boundary at $r=h$ is crossed, it is impossible to return or communicate with the exterior.
-This Ref.
-Ovalle (2024) examines whether an alternative structure, beyond a pointlike mass, can give rise to the Schwarzschild black hole region for $r\geq h$.
+In this line of research, remarkably, the idea that an interior region, rather than a pointlike mass, can generate a Schwarzschild black hole exterior has recently gained attention Ovalle (2024); Maeda (2025).
+This configuration is obtained by matching an interior solution to the Schwarzschild exterior at the event horizon, without introducing a lightlike thin shell.
+Reference Ovalle (2024) examines whether an alternative structure, beyond a pointlike mass, can give rise to the Schwarzschild black hole region for $r\geq h$.
 In particular, the authors aim to describe an interior configuration associated with this region while keeping the exterior Schwarzschild geometry unchanged and ensuring that tidal forces remain finite everywhere.
 In connection with the latter, the authors propose that the internal region may contain either a central singularity or a regular core endowed with a potentially unstable inner horizon.
-It is well known that a wide class of black hole solutions studied in the literature are supported by matter sources that inevitably lead to a central singularity, for example BH cosmic-void density profiles Lustosa et al.
-(2025), BH dark-matter with dark matter energy profile Xu et al.
-(2018), and hairy BHs Ovalle et al.
+Reference Maeda (2025) constructs a so-called fake Schwarzschild solution, observationally indistinguishable from the Schwarzschild black hole, by matching a dynamical spherically symmetric interior containing matter to the Schwarzschild exterior at the event horizon without introducing a lightlike thin shell.
+It is shown that this configuration can be obtained from a perfect-fluid solution satisfying a linear equation of state of the type considered in Ref.
+Maeda (2023), analytically continued inside the horizon.
+It is well known that a wide class of black hole solutions studied in the literature are supported by matter sources that inevitably lead to a central singularity, just to mention a few examples: black holes with cosmic-void density profiles Lustosa et al.
+(2025), black holes surrounded by dark matter with dark energy profiles Xu et al.
+(2018), hairy black holes Ovalle et al.
+(2021), and quintessential black holes Kiselev (2003); Estrada et al.
 (2025).
-Electrically charged BH, exemplified by the Reissner–Nordström solution, also arise from electromagnetic sources in the energy–momentum tensor and are characterized by the presence of an inner horizon.
+Furthermore, electrically charged BH, exemplified by the Reissner–Nordström solution, also arise from electromagnetic sources in the energy–momentum tensor and are characterized by the presence of an inner horizon.
 Motivated by the above considerations, and in analogy with Ref.

### v3

@@
 Abstract
-Regular black holes are often used to address singularities, but they typically involve a potentially unstable de Sitter core and an internal horizon that breaks predictability.
-Integrable singularities (IS) have recently gained attention because they avoid both issues and exhibit finite tidal forces, allowing nondestructive radial infall.
-First, we present a new BH solution sourced by a string fluid (FS) that exhibits an IS.
-Motivated by the divergence of the conserved energy in the cloud of strings (CS) model, we introduce an energy density profile based on the screening of the CS energy density within an FS framework, thereby yielding finite conserved energy.
-On the other hand, the idea that an interior region, rather than a pointlike mass, can generate a Schwarzschild black hole exterior region has recently gained attention Ovalle (2024); Maeda (2025).
-This is obtained by matching an interior solution to the Schwarzschild exterior at the event horizon.
-Secondly, motivated by the variety of singular BH solutions in the literature, we establish the conditions that an interior region with an IS must satisfy in order to represent the interior of a generic exterior BH solution, with Schwarzschild being only a particular case of the latter.
-We derive the junction conditions (JC) between the interior and exterior regions, showing that they lead to temperature continuity at the interface, while discontinuities in tangential pressure lead to phase transitions.
-We propose that the nature of the interior region is described by CS and FS, while the exterior corresponds to Reissner–Nordström.
+BHs with integrable singularities (IS) have gained attention because, unlike usual regular BHs, they avoid a potentially unstable de Sitter core and the presence of an internal horizon that breaks predictability, while exhibiting finite tidal forces that allow nondestructive radial infall.
+First, we present a new BH solution sourced by a string fluid (FS) exhibiting an IS.
+We introduce an energy density profile obtained by screening the cloud of strings (CS) density within an FS framework, leading to finite conserved energy, unlike the CS case where it diverges.
+On the other hand, the idea that an interior region, rather than a pointlike mass, can generate a Schwarzschild exterior BH has recently gained attention Ovalle (2024); Maeda (2025).
+This is achieved by matching an interior region to the Schwarzschild exterior at the event horizon.
+Motivated by the variety of singular BH solutions in the literature, we establish the conditions that an interior region with an IS must satisfy to represent the interior of a generic exterior BH solution, with Schwarzschild as only a particular case.
+We derive the junction conditions between the interior and exterior regions, showing that they ensure temperature continuity at the interface, while discontinuities in tangential pressure signal phase transitions.
+We examine an interior described by CS and FS, matched to a Reissner–Nordström exterior.
 I Introduction

### v4

@@
 Abstract
-BHs with integrable singularities (IS) have gained attention because, unlike usual regular BHs, they avoid a potentially unstable de Sitter core and the presence of an internal horizon that breaks predictability, while exhibiting finite tidal forces that allow nondestructive radial infall.
-First, we present a new BH solution sourced by a string fluid (FS) exhibiting an IS.
-We introduce an energy density profile obtained by screening the cloud of strings (CS) density within an FS framework, leading to finite conserved energy, unlike the CS case where it diverges.
-On the other hand, the idea that an interior region, rather than a pointlike mass, can generate a Schwarzschild exterior BH has recently gained attention Ovalle (2024); Maeda (2025).
-This is achieved by matching an interior region to the Schwarzschild exterior at the event horizon.
-Motivated by the variety of singular BH solutions in the literature, we establish the conditions that an interior region with an IS must satisfy to represent the interior of a generic exterior BH solution, with Schwarzschild as only a particular case.
-We derive the junction conditions between the interior and exterior regions, showing that they ensure temperature continuity at the interface, while discontinuities in tangential pressure signal phase transitions.
-We examine an interior described by CS and FS, matched to a Reissner–Nordström exterior.
+Black Holes (BHs) with integrable singularities (IS) have attracted attention because, unlike usual regular BHs, they avoid a potentially unstable de Sitter core and the presence of an internal horizon that breaks predictability, while exhibiting finite tidal forces that allow nondestructive radial infall.
+On the other hand, the idea that an interior region, rather than a pointlike mass, can generate a Schwarzschild exterior BH region has recently gained attention Ovalle (2024); Maeda (2025).
+In this work, motivated by the existence of an inner horizon and a non integrable singularity in the Reissner Nordström solution (RN), we address the possibility that the RN BH possesses an interior region with an IS and no inner horizon, instead of a pointlike mass.
+We propose that the nature of this interior region is given by a cloud of strings (CS) with IS and by a new fluid of strings (FS) with IS, determining the physical properties that both must satisfy in order to support such a configuration.
+To this end, we establish the matching conditions between these interior regions with IS and the RN exterior BH solution.
+This leads to the requirement that the continuity of temperature must be satisfied at the interface, while discontinuities in the tangential pressure signal phase transitions.
+Thus, for both interior regions, we determine the conditions relating them to the RN exterior region that must be satisfied in order to physically represent the interior region with an IS while fulfilling the matching conditions.
+Furthermore, we identify the conditions under which phase transitions occur or do not occur.
 I Introduction
@@
 The dynamics of a single 1-brane string can be extended to an ensemble Yang et al.
-(2023), forming a cloud of strings.
+(2023), forming a cloud of strings (CS).
 Embedding such a cloud of strings modifies the spacetime geometry, producing black hole solutions surrounded by cloud of strings Letelier (1979), and several recent studies have analyzed black holes interacting with string clouds Zafar et al.
@@
 (2025).
-A generalization of this model includes pressure or tension in the energy–momentum tensor, giving it a structure similar to a perfect fluid, i.e., a fluid of strings Letelier (1983).
-In this connection, recently an equation of state $\rho(r)=\alpha(r)p(r)$ was proposed Nunes dos Santos (2025) for a fluid of strings, leading to a regular black hole solution.
-Following Ref.
+A generalization of this model includes pressure or tension in the energy–momentum tensor, giving it a structure similar to a perfect fluid, i.e., a fluid of strings (FS) Letelier (1983).
+In this connection, recently an equation of state $\rho(r)=\alpha(r)p(r)$ was proposed Nunes dos Santos (2025) for a fluid of strings, leading to a regular black hole model.
+Following the definition of the conserved energy associated with a timelike Killing vector given in Ref.
 Aoki et al.
-(2021), where the volume integral $\int_{0}^{\infty}4\pi r^{2}T^{0}{}_{0},dr$ is interpreted as the conserved energy associated with a timelike Killing vector, we note that this integral diverges for the standard string-cloud energy density.
-Motivated by this, as the first problem in this work we introduce a modified version of the string-cloud energy density within the fluid-of-strings framework.
+(2021), we note that this conserved energy diverges for the standard cloud of strings energy density.
+Motivated by this, we first introduce in this work a modified version of the string-cloud energy density within the fluid-of-strings framework.
 This modification can be interpreted as a geometrical screening of the original distribution.
-As a result, we show that the volume integral converges to a finite value equal to $M$, which can be interpreted as the conserved energy charge.
-In this way, we provide a new fluid-of-strings black hole, constructed through the equation of state $\rho(r)=\alpha(r)p(r)$.
-Furthermore, we will show that the Ricci scalar near the origin behaves in a way consistent with an integrable singularity, and that the new solution is also free of a potentially unstable inner horizon.
+As a result, this leads to the convergence of the conserved energy charge.
+In this way, we first provide a new fluid-of-strings black hole model, constructed through the equation of state $\rho(r)=\alpha(r)p(r)$.
+Furthermore, we show that the Ricci scalar near the origin behaves in a way consistent with an integrable singularity, and that the new FS model is also free of a potentially unstable inner horizon.
 On the other hand, the Schwarzschild solution, which according to Birkhoff’s theorem Hawking and Ellis (2023) is the unique spherically symmetric vacuum solution of Einstein’s equations, possesses a strong central singularity.
-In this paragraph, we point out some ideas discussed in Ref.
-Ovalle (2024).
-The Schwarzschild black hole features an event horizon, which can be viewed as a membrane of radius $h$ separating two causally disconnected space-time regions: an interior and an exterior one.
+As pointed out in reference Ovalle (2024), the Schwarzschild black hole features an event horizon, which can be viewed as a membrane of radius $h$ separating two causally disconnected space-time regions: an interior and an exterior one.
 The interior region, defined by $r<h$, contains the central singularity shielded by the event horizon.
@@
 In particular, the authors aim to describe an interior configuration associated with this region while keeping the exterior Schwarzschild geometry unchanged and ensuring that tidal forces remain finite everywhere.
-In connection with the latter, the authors propose that the internal region may contain either a central singularity or a regular core endowed with a potentially unstable inner horizon.
-Reference Maeda (2025) constructs a so-called fake Schwarzschild solution, observationally indistinguishable from the Schwarzschild black hole, by matching a dynamical spherically symmetric interior containing matter to the Schwarzschild exterior at the event horizon without introducing a lightlike thin shell.
+In connection with the latter, the authors propose that the internal region may contain either a central singularity or a regular core.
+Reference Maeda (2025) constructs a so-called fake Schwarzschild model, observationally indistinguishable from the Schwarzschild black hole, by matching a dynamical spherically symmetric interior containing matter to the Schwarzschild exterior at the event horizon without introducing a lightlike thin shell.
 It is shown that this configuration can be obtained from a perfect-fluid solution satisfying a linear equation of state of the type considered in Ref.
 Maeda (2023), analytically continued inside the horizon.
-It is well known that a wide class of black hole solutions studied in the literature are supported by matter sources that inevitably lead to a central singularity, just to mention a few examples: black holes with cosmic-void density profiles Lustosa et al.
-(2025), black holes surrounded by dark matter with dark energy profiles Xu et al.
-(2018), hairy black holes Ovalle et al.
-(2021), and quintessential black holes Kiselev (2003); Estrada et al.
+Reissner Nordström solution represents an electrically charged BH, arising from electromagnetic sources in the energy momentum tensor and characterized by the presence of an inner horizon and a central singularity.
+Motivated by the considerations described in this introduction, and by the existence of an inner horizon and a non-integrable singularity in the Reissner–Nordström solution (RN), in this work, we study the physical and thermodynamic properties of a scenario in which the RN BH possesses an interior region with an integrable singularity and no inner horizon, instead of a pointlike mass.
+We investigate the possibility that the nature of the matter sources capable of supporting such an internal geometry with an integrable singularity corresponds to those associated with clouds of strings and with the new string fluid introduced in this work.
+To carry out this work, we will first define the conditions that a general internal geometry must satisfy in order to represent an integrable singularity, as well as the Israel Darmois Israel (1966) junction conditions between a general internal region and a general external BH geometry.
+We will then examine the connection between the junction conditions and the thermodynamic properties of BHs, and explore their relation to the existence of phase transitions.
+In this context, we will show that satisfying the junction conditions leads to the requirement that the continuity of temperature must hold at the interface, while discontinuities in the tangential pressure signal phase transitions.
+It is worth mentioning that these general conditions could also be of physical interest for future studies of other BH exterior regions, instead of RN, that possess a central singularity, for example BHs with cosmic void density profiles Lustosa et al.
+(2025), BHs surrounded by dark matter Xu et al.
+(2018), hairy BHs Ovalle et al.
+(2021), and quintessential BHs Kiselev (2003); Estrada et al.
 (2025).
-Furthermore, electrically charged BH, exemplified by the Reissner–Nordström solution, also arise from electromagnetic sources in the energy–momentum tensor and are characterized by the presence of an inner horizon.
-Motivated by the above considerations, and in analogy with Ref.
-Ovalle (2024), we address as a second problem the existence of an internal geometry inside the event horizon of a generic black hole spacetime, keeping the external geometry completely general, with the Schwarzschild solution appearing only as a particular case.
-We focus on scenarios in which the internal structure is endowed with an integrable singularity, in order to avoid the emergence of a potentially unstable de Sitter core and the presence of an inner horizon, since predictability breaks down after crossing the inner horizon.
-We investigate the possible nature of matter sources capable of supporting such an internal geometry with an integrable singularity, specifically those associated with clouds of strings and with the new string fluid introduced in this work.
-To do this, we define the conditions that the internal geometry must satisfy in order to represent an integrable singularity, as well as the junction conditions between generic internal and external black hole geometries.
-Furthermore, we examine the connection between the Israel–Darmois Israel (1966) junction conditions and black hole thermodynamic properties, and explore their relation to the existence of phase transitions.
-As a test case, we study the Reissner–Nordström solution as an internal geometry, considering clouds of strings and the new string fluid as the first and second examples, respectively.
-II A new Fluid of strings black hole solution with an integrable singularity
-In this section, we first present a new string fluid solution, showing that it can represent a black hole solution over the entire domain $r\in[0,\infty)$.Below, in line with this work, we characterize it as the interior region of a Reissner–Nordström black hole.
+In this way, applying the general conditions described above, in this work we focus on the physical and thermodynamic properties of the mentioned configuration where the exterior geometry is given by the Reissner Nordström BH, while the interior geometry is described by a cloud and by a new fluid of strings with integrable singularities.
+Specifically, we will determine the physical properties that both the cloud of strings (CS) and the new fluid of strings (FS) must satisfy in order to represent the interior of an exterior RN BH region.
+Using the general analysis described above, we will determine the relations between the parameter spaces of both interior regions and the RN exterior BH for which the matching conditions are satisfied and, as a consequence, the temperature remains continuous at the interface.
+In addition, for both interior regions we will determine the conditions relating them to the RN exterior region that must be satisfied for phase transitions to occur or not occur.
+II A new Fluid of strings black hole model with an integrable singularity
+In this section, we first present a new string fluid model, showing that it can represent a black hole model over the entire domain $r\in[0,\infty)$.
+Below, in line with this work, we characterize it as the interior region of a Reissner–Nordström black hole.
 Reference Letelier (1983) proposes a generalization of the cloud of strings model by incorporating pressure or tension into the energy–momentum tensor.
-This model is the so-called fluid of strings (fs).
+This model is the so-called fluid of strings.
 See Appendix B for further details.
@@
 Nunes dos Santos (2025), an equation of state was proposed in which the energy density and the pressure are related through the form $\rho(r)^{(fs)}=\alpha(r)\cdot p(r)^{(fs)}$.
-In this reference, an equation of state with $\alpha(r)$ was proposed such that the solution of Einstein’s equations represents a regular string-fluid geometry at the origin.
+In this reference, an equation of state with $\alpha(r)$ was proposed such that the model of Einstein’s equations represents a regular string-fluid geometry at the origin.
 Within this framework, the energy–momentum tensor takes the form
@@
 As indicated in Ref.
-Nunes dos Santos (2025), within this framework the Einstein equations lead to a solution of the form:
+Nunes dos Santos (2025), within this framework the Einstein equations lead to a geometry of the form:
 $f(r)_{(fs)}=1+\frac{c_{2}}{r}+\frac{c_{1}}{r}\,I(r),$ (3)
@@
 1 .
-Consequently, our solution lacks the presence of an unstable de Sitter core, unlike other black hole solutions sourced by matter.
+Consequently, our model lacks the presence of an unstable de Sitter core, unlike other black hole solutions sourced by matter.
 Figure 1: $f(r)_{(fs)}$ for $b=1,M=0.8,0.9,1$. We observe the absence of a potentially unstable inner horizon.
@@
 For both reasons discussed above, we can conclude that there is no presence of a potentially unstable inner horizon.
-In this way, in this section we have presented a new string fluid solution that represents a geometry with an integrable singularity and the absence of a potentially unstable inner horizon.
-As mentioned above, in this section we analyze a black hole solution over the entire domain $r\in[0,\infty)$.
+In this way, in this section we have presented a new string fluid model that represents a geometry with an integrable singularity and the absence of a potentially unstable inner horizon.
+As mentioned above, in this section we analyze a black hole model over the entire domain $r\in[0,\infty)$.
 Below, in line with this work, we characterize it as the interior region of a Reissner–Nordström black hole.
-III Black Hole Solutions with an Interior Geometry Featuring an Integrable Singularity
+III Black Hole models with an Interior Geometry Featuring an Integrable Singularity
 In Fig.
@@
 $\frac{\left(r^{2}f^{\prime}\right)^{\prime}}{2r^{2}}+\Lambda=$ $p_{\theta}$ (18)
-where we observe that $\rho=-p_{r}$. The conservation equation reads:
+where we observe that $\rho=-p_{r}$.
 Thus, we can define the effective interior energy-momentum tensor $(\bar{T}^{\mu}_{\ \nu})^{(\text{in})}$, for $\Lambda=-3/l^{2}$, where $l$ denotes the AdS radius.
@@
 Since this region represents the exterior of the black hole, the signature of the metric tensor (25) is ,+,+,+-,+,+,+, and therefore $f_{E}(r)\geq 0$ in this region.
-We consider the case without a cosmological constant, with an energy–momentum tensor of the form
+We consider the case without a cosmological constant in the external region, with an energy–momentum tensor of the form
 $(T^{\mu}_{\ \nu})^{(E)}=\mathrm{diag}\left(-\rho^{(E)},p_{r}^{(E)},p_{\theta}^{(E)},p_{\theta}^{(E)}\right)$ (26)
@@
 IV INTERIOR REGION AS AN INTEGRABLE SINGULARITY
-In this section, we state the properties that a generic interior solution must satisfy in order to represent an integrable singularity near the origin.
-In the following sections, we analyze the specific cases in which the interior regions correspond, respectively, to a string cloud and to the new string-fluid solution provided in Section II.
+In this section, we state the properties that a generic interior geometry must satisfy in order to represent an integrable singularity near the origin.
+In the following sections, we analyze the specific cases in which the interior regions correspond, respectively, to a string cloud and to the new string-fluid model provided in Section II.
 The trace of our efective interior energy–momentum tensor (19) is given by $\bar{T}^{(in)}=-2\bar{\rho}+2\bar{p}_{\theta}$. Replacing Eqs. (23) and (24), we obtain
@@
 As mentioned above, the absence of zeros in this interval implies the absence of a potentially unstable inner horizon.
-Although the Ricci scalar has a singularity $\sim r^{-2}$ at the origin, the equations of motion (28) are integrable, being free of singularities, and lead to a finite solution $f(r)$ at the origin.
+Although the Ricci scalar has a singularity $\sim r^{-2}$ at the origin, the equations of motion (28) are integrable, being free of singularities, and lead to a finite geometry $f(r)$ at the origin.
 This situation can be referred to as an integrable singularity.
@@
 In this way, $df/dr>0$ for $r\in[0,h]$, with the temperature $T=\left.df/dr\right|_{h^{-}}=\left.df_{E}/dr\right|_{h^{+}}$.
+It is worth mentioning that this condition leads to the absence of an inner horizon.
 V Junction conditions
@@
     As we will see below, this condition is not mandatory, unlike the two previous ones.
-More specifically, this condition represents a consequence of the discontinuity (or continuity) of the tangential pressure in black hole thermodynamics.We obtain the interior temperature from Eq.
+More specifically, this condition represents a consequence of the discontinuity (or continuity) of the tangential pressure in black hole thermodynamics.
+We obtain the interior temperature from Eq.
 (24).
@@
 However, the criteria introduced above must be tested in explicit black hole geometries.
-In this spirit, we analyze two illustrative examples.
+In this spirit, motivated by the existence of an inner horizon and a non-integrable singularity in the Reissner–Nordström solution (RN), we analyze two illustrative examples.
 In the present section, we study a Reissner–Nordström exterior with a cloud of strings as the interior region.
-In the following section, we consider a Reissner–Nordström exterior with the new string-fluid solution provided in Section  II.
+In the following section, we consider a Reissner–Nordström exterior with the new string-fluid model provided in Section  II.
 Remarks on the Reissner–Nordström (RN) Black Hole:
@@
 $h=M+\sqrt{M^{2}-Q^{2}}$ (41)
-Thus, we note that it must be satisfied that $M>Q\Rightarrow M^{2}-Q^{2}>0$.
+Thus, we note that it must be satisfied that $M>Q$ for $M,Q>0$, $\Rightarrow M^{2}-Q^{2}>0$.
 As we have mentioned, in our setup, the exterior geometry extends over $r\in[h,r_{b}]$, with the boundary corresponding to the radial coordinate approaching infinity, since the solution is asymptotically flat.
@@
 $\rho=\frac{a}{r^{2}}$ (46)
-Thus, from equations (20),(21), (22), we note that the interior energy density and effective pressures take the following form
+Thus, from equations (20),(21), (22), we note that the interior effective energy density and effective pressures take the following form
 $\bar{\rho}=-\bar{p}_{r}=\frac{a}{r^{2}}-\frac{3}{l^{2}}\,\,,\,\,\bar{p}_{\theta}=\frac{3}{l^{2}}$ (47)
@@
 This result is further reinforced by the fact that, in this work, we are interested in an interior geometry associated with an integrable singularity, for which the Ricci tensor is such that the equations of motion remain integrable.
-In order to test condition IV 2 for the interior region to be represented as an integrable singularity, we note that $f(r)$ is an increasing function, starting from its minimum value f(r=0)=1a<0f(r=0)=1-a<0, with $a>1$, and reaching $f(h)=0$.
+In order to test conditions IV and 2 for the interior region to be represented as an integrable singularity without an inner horizon, we note that $f(r)$ is an increasing function, starting from its minimum value f(r=0)=1a<0f(r=0)=1-a<0, with $a>1$, and reaching $f(h)=0$.
 Indeed, $df/dr=2r/l^{2}>0$ for $r\in[0,h]$, and therefore the function is monotonically increasing and negative throughout this interval.
@@
     $C_{\text{in}}-C_{E}\sim\bar{p}_{\theta}(h)-p_{\theta}^{(E)}(h)=-\Lambda-\frac{Q^{2}}{h^{4}}=\frac{3M\left(M+\sqrt{M^{2}-Q^{2}}\right)-4Q^{2}}{\left(M+\sqrt{M^{2}-Q^{2}}\right)^{4}}.$ (52)
-    Testing the denominator of the last equation for zero leads us to:
+    Testing the numerator of the last equation for zero leads us to:
     $Q^{2}\left(\frac{16Q^{2}}{9M^{2}}-\frac{5}{3}\right)=0\Rightarrow Q_{c}^{2}=\frac{15}{16}M^{2}$ (53)
@@
 VII Example 2: A Reissner–Nordström Exterior with a new fluid of Strings as the Interior Region with an Integrable Singularity
-In Section II, a new string fluid solution was introduced.
+In Section II, a new string fluid model was introduced.
 There, it was presented under the assumption that it extends over the entire spacetime, $r\in[0,\infty)$.
-In the present section, we analyze instead the case in which the solution of Section II, in the presence of a negative cosmological constant, is defined only in the interval $r\in[0,h]$, i.e., it can represent the interior region of the Reissner–Nordström solution, as depicted in Scheme 2.
+In the present section, we analyze instead the case in which the model of Section II, in the presence of a negative cosmological constant, is defined only in the interval $r\in[0,h]$, i.e., it can represent the interior region of the Reissner–Nordström solution, as depicted in Scheme 2.
 It is worth mentioning that, in Section VI, we have already highlighted the main features of the Reissner–Nordström solution that are relevant to this work.
-VII.1 Representation of the New Fluid of Strings solution as an Interior Region with an Integrable Singularity
+VII.1 Representation of the New Fluid of Strings model as an Interior Region with an Integrable Singularity
 The effective interior energy–momentum tensor (19), obtained from Eqs. (20), (21), and (22), (5) and (11), takes the following form
@@
 Consequently, condition IV 1 is satisfied, allowing this new string fluid model to consistently represent the interior region as an integrable singularity.
-The solution is given by:
+The geometry is given by:
 $f(r)=1-\frac{2M}{r}\left(1-\exp(-r/b)\right)+\frac{r^{2}}{l^{2}}$ (57)
 which behaves as $f\mid_{r\sim 0}\sim 1-\frac{2M}{b}+\frac{r^{2}}{l^{2}}$ near the origin, with $\frac{2M}{b}>1$ and $|1-\frac{2M}{b}|$ sufficiently large compared to the term $\frac{r^{2}}{l^{2}}$ in the vicinity of the origin, so as to avoid a zero of the function $f(r)$ near the origin.
-In order to test condition IV 2 for the interior region to be represented as an integrable singularity, we note that $f(r)$ is an increasing function, starting from its minimum value $f(r=0)=1-\frac{2M}{b}<0$, with $\frac{2M}{b}>1$, and, as we shall see below, reaching $f(h)=0$.
+In order to test condition IV 2 for the interior region to be represented as an integrable singularity without an inner horizon, we note that $f(r)$ is an increasing function, starting from its minimum value $f(r=0)=1-\frac{2M}{b}<0$, with $\frac{2M}{b}>1$, and, as we shall see below, reaching $f(h)=0$.
 Indeed, $df/dr=\frac{2M}{r^{2}}\left(1-\exp(-r/b)\left(1+\frac{r}{b}\right)\right)+\frac{2r}{l^{2}}>0$ for $r\in[0,h]$.
@@
 VIII Discussion and Summary
-In Section II, we present a new fluid-of-strings (fs) black hole solution defined over the entire domain $r\in[0,\infty)$ and characterized by the equation of state $\rho^{(fs)}(r)$ and $p^{(fs)}(r)$.
-Following Ref.
-Aoki et al.
-(2021), where the integral $\int_{0}^{\infty}4\pi r^{2}T^{0}{}_{0},dr$ is interpreted as the conserved energy associated with a timelike Killing vector, we note that this integral diverges for the standard string-cloud energy density.
-Motivated by this behavior, we introduce a modified energy density for the fluid of strings, which can be interpreted as a geometrical screening of the original string-cloud distribution.
-As a result, the aforementioned integral converges to a finite value equal to $M$, which can be interpreted as the conserved energy charge.
-Although the Ricci scalar diverges near the origin as $R\sim r^{-2}$, these divergences are integrable, and the spacetime remains free of curvature singularities, allowing the solution to be classified as an integrable singularity.
-Finally, we show that the solution is free of a potentially unstable de Sitter core and of an inner horizon, whose presence is associated with a breakdown of causality.
-Motivated by the fact that several black hole solutions sourced by matter fields leading to a destructive central singularity have attracted considerable interest in the literature, and in analogy with Ref.
-Ovalle (2024), we have addressed, as a second problem, the existence of an internal geometry inside the event horizon of a generic black hole spacetime.
-Within this framework, the external geometry is kept completely general, with the Schwarzschild solution appearing only as a particular case.
-We have focused on scenarios in which the internal structure is endowed with an integrable singularity, in order to avoid the emergence of a potentially unstable de Sitter core and the presence of an inner horizon, since predictability breaks down after crossing the inner horizon.
-We have investigated the nature of the matter sources capable of supporting such an internal geometry with an integrable singularity, specifically those associated with clouds of strings and with the new string fluid introduced in this work.
-To this end, in Section IV we have defined the conditions that the internal geometry must satisfy in order to represent an integrable singularity.
-In Section V we have established the junction conditions between the internal and generic external black hole geometries.
-We have shown that the second fundamental form of the Israel–Darmois junction conditions Israel (1966) leads to the continuity of the temperature across the interface between the internal and external regions.
-Furthermore, from the analysis of the junction conditions, we have determined that a discontinuity in the tangential pressure implies the existence of a phase transition at the location of the event horizon, and vice versa.
-In Sections VI and VI, we studied the cases in which the internal geometry of the Reissner–Nordström spacetime is described, respectively, by a string cloud and by our newly proposed string fluid geometry.
+In this work, motivated by the existence of an inner horizon and a non-integrable singularity in the Reissner–Nordström solution (RN), we study the physical and thermodynamic properties of a scenario in which the RN black hole possesses an interior region with an integrable singularity and no inner horizon, instead of a pointlike mass.
+We have investigated the possibility that the nature of the matter sources capable of supporting such an internal geometry with an integrable singularity (IS) corresponds to clouds of strings (CS) and fluids of strings (FS).
+In this connection, motivated by the fact that the conserved charge energy in the CS case diverges, we have proposed a model for the internal region based on an FS in which the energy density, which can be interpreted as a geometrical screening of the original CS distribution, leads to a finite conserved charge energy.
+We have shown that both the CS and FS models used in this work can represent an interior region with an IS and that they lack an inner horizon where predictability would be lost.
+To carry out this work, in Section IV we have determined the conditions that a general internal geometry must satisfy in order to represent an integrable singularity.
+In Section V we have defined the Israel–Darmois Israel (1966) junction conditions between a general internal region and a general external BH geometry.
+In this context, we have shown that satisfying the junction conditions requires the continuity of temperature at the interface, while discontinuities in the tangential pressure signal the presence of phase transitions.
+It is worth mentioning that these general conditions are also of interest for future studies of other BH exterior regions that possess a central singularity instead of the RN one, for example BHs with cosmic void density profiles Lustosa et al.
+(2025), BHs surrounded by dark matter Xu et al.
+(2018), hairy BHs Ovalle et al.
+(2021), and quintessential BHs Kiselev (2003); Estrada et al.
+(2025).
+Using the framework developed in the previous paragraph, in Sections VI and VII we studied the cases in which the internal geometry of the Reissner–Nordström spacetime is described, respectively, by a string cloud and by our newly proposed string fluid model.
 In both scenarios, the parameters of the corresponding internal geometries satisfy the conditions listed in Section IV, ensuring the presence of an integrable singularity near the origin, the absence of a potentially unstable de Sitter core, and the absence of an internal horizon that could lead to a breakdown of causality.
@@
 In addition, we identified a critical value of the electric charge for which no phase transition occurs in the location of the interface between the two regions.
-In Section VI, we derived the conditions that the AdS radius, the electric charge, the screening parameter $b$, and the event horizon must satisfy in order to fulfill both Israel junction conditions at the interface between the external RN region and the internal region described by the new string fluid.
+In Section VII, we derived the conditions that the AdS radius, the electric charge, the screening parameter $b$, and the event horizon must satisfy in order to fulfill both Israel junction conditions at the interface between the external RN region and the internal region described by the new string fluid.
 Moreover, we numerically determined the critical value of the screening parameter for which no phase transition occurs in the location of the interface between the two regions.
-Appendix A A brief revision of the cloud of string solution
+Appendix A A brief revision of the cloud of string model
 In this appendix, we review the necessary tools for our study of the so-called cloud of strings from reference Letelier (1979).
@@
 where $\rho_{cs}$ can be interpreted as the proper rest energy density of the string cloud.
-Appendix B Some general aspects of the fluid of strings solution
+Appendix B Some general aspects of the fluid of strings model
 As previously discussed, the original cloud of strings framework was subsequently extended in Ref.

### v5

@@
-
+Reissner Nordström black holes with integrable singularity interiors supported by string distributions
 Abstract
-Black Holes (BHs) with integrable singularities (IS) have attracted attention because, unlike usual regular BHs, they avoid a potentially unstable de Sitter core and the presence of an internal horizon that breaks predictability, while exhibiting finite tidal forces that allow nondestructive radial infall.
-On the other hand, the idea that an interior region, rather than a pointlike mass, can generate a Schwarzschild exterior BH region has recently gained attention Ovalle (2024); Maeda (2025).
-In this work, motivated by the existence of an inner horizon and a non integrable singularity in the Reissner Nordström solution (RN), we address the possibility that the RN BH possesses an interior region with an IS and no inner horizon, instead of a pointlike mass.
-We propose that the nature of this interior region is given by a cloud of strings (CS) with IS and by a new fluid of strings (FS) with IS, determining the physical properties that both must satisfy in order to support such a configuration.
-To this end, we establish the matching conditions between these interior regions with IS and the RN exterior BH solution.
-This leads to the requirement that the continuity of temperature must be satisfied at the interface, while discontinuities in the tangential pressure signal phase transitions.
-Thus, for both interior regions, we determine the conditions relating them to the RN exterior region that must be satisfied in order to physically represent the interior region with an IS while fulfilling the matching conditions.
-Furthermore, we identify the conditions under which phase transitions occur or do not occur.
+The Reissner Nordström (RN) black hole is characterized by two well known pathologies: a central singularity and an inner horizon associated with instabilities and a potential loss of predictability.
+In this work, we show that the RN exterior geometry can arise from an interior spacetime containing an integrable singularity but no inner horizon.
+In this scenario, tidal forces remain finite near the origin, allowing nondestructive radial infall, while the conventional description in terms of a pointlike mass is replaced by an extended matter distribution.
+To illustrate this possibility, we provide explicit realizations of such an interior region based on string distributions, namely a cloud of strings (CS) and a newly defined fluid of strings (FS).
+While the standard cloud of strings model leads to a divergence in the conserved energy associated with timelike Killing vectors, the proposed FS model can be interpreted as a geometrically screened version of the string cloud distribution and admits configurations that, when extended to infinity, describe black holes with finite conserved energy.
+Physical consistency between the interior region and the RN exterior geometry requires the continuity of temperature across the interface, implying thermal equilibrium between the two regions, while discontinuities in the tangential pressure can signal gravitational phase transitions.
+These results determine the physical conditions under which string based interior distributions can consistently generate the RN exterior geometry and clarify the circumstances under which phase transitions at the event horizon may arise.
 I Introduction
@@
 (2016) has firmly positioned black holes among the most compelling phenomena in gravitational physics.
-In general, black hole solutions are characterized by the presence of a central singularity, where the spacetime description breaks down, signaled by the divergence of the curvature invariants.
-To address this problem, the standard approach involves specifying a particular form of matter within the energy-momentum tensor, which gives rise to regular black holes (RBHs).
-It is well established that RBH solutions typically include an inner horizon, which can also lead to the presence of a de Sitter–like core near the origin.
-In this context, several studies have indicated that the inner horizon is inherently unstable Poisson and Israel (1990); Brown et al.
+Black hole solutions are typically associated with the presence of a central singularity, where the geometric description of spacetime breaks down due to the divergence of curvature invariants.
+A widely studied approach to address this issue is provided by regular black holes (RBHs), in which the central singularity is usually replaced by a de Sitter core.
+However, these models generally possess an inner horizon, whose presence leads to well-known dynamical problems.
+In particular, the phenomenon of mass inflation can generate severe instabilities at this horizon Poisson and Israel (1990); Brown et al.
 (2011).
-Along the same lines, a more recent study, reference Carballo-Rubio et al.
-(2023), argues that instabilities at the cores of regular black holes associated with the inner horizon are inevitable, since the mass inflation phenomenon represents a fundamental feature of RBHs with astrophysical significance.
-On the other hand, it is also worth noting that, as claimed in reference Ovalle (2023), predictability breaks down after crossing the inner horizon.
-As an alternative to RBH solutions, Ref.
-Lukash and Strokov (2013) proposes a gravitational scenario in which a particle can reach the radial origin without being destroyed.
-Rather than a potentially unstable de Sitter core associated with an inner horizon, this model exhibits an integrable singularity at the origin.
-Integrable singularities are characterized by a divergence of the Ricci scalar while their volume integrals remain finite near the center.
-The elimination of the inner horizon thus prevents possible instabilities.
-Furthermore, recent studies in Ref.
-Estrada et al.
-(2024) show that a central integrable singularity allows for finite tidal forces, extendable radial geodesics, and a weak singularity, as formalized by Tipler’s extension theorem Nolan (2000).
-This ensures that an object approaching the central singularity would not undergo spaghettification.
-In addition, Ref.
-Casadio et al.
-(2023) offers a quantum perspective on the existence of such integrable singularities.
-For other recent applications, see Refs.
-Ovalle (2023); Estrada (2025); Arrechea et al.
+More recent studies have reinforced this conclusion, indicating that instabilities associated with the inner horizon appear to be an unavoidable feature of astrophysically relevant RBHs Carballo-Rubio et al.
 (2023).
-On the other hand, String theory posits that the fundamental constituents of the universe are one-dimensional objects called strings Younesizadeh and Jokar (2022).
-The dynamics of a single 1-brane string can be extended to an ensemble Yang et al.
-(2023), forming a cloud of strings (CS).
-Embedding such a cloud of strings modifies the spacetime geometry, producing black hole solutions surrounded by cloud of strings Letelier (1979), and several recent studies have analyzed black holes interacting with string clouds Zafar et al.
+Moreover, it has been argued that the existence of such an inner horizon may compromise the predictability of spacetime beyond it Ovalle (2023).
+An alternative route that has recently attracted attention consists of replacing the regular core with an integrable singularity (IS) Lukash and Strokov (2013).
+In this type of configuration, curvature invariants may diverge at the origin, while their volume integrals remain finite.
+An important consequence of this property is that tidal forces experienced by infalling observers can remain finite, suggesting that objects approaching the singularity radially may not necessarily be destroyed.
+Recent studies, such as Estrada et al.
+(2024), have shown that these singularities can be weak according to Tipler’s classification Nolan (2000) and may allow the extension of radial geodesics through the origin .
+Furthermore, possible quantum interpretations of integrable singularities have also been explored Casadio et al.
+(2023).
+Additional applications of this framework can be found in Ovalle (2023); Estrada (2025); Arrechea et al.
+(2025).
+On the other hand, string theory suggests that the fundamental constituents of matter may be described as one-dimensional objects known as strings Younesizadeh and Jokar (2022).
+In an effective gravitational framework, a macroscopic ensemble of such objects can be modeled as a cloud of strings, whose energy–momentum tensor modifies the surrounding spacetime geometry Letelier (1979).
+This type of source has been widely used to study black hole models surrounded by string distributions Zafar et al.
 (2025); Javed et al.
@@
 (2025).
-A generalization of this model includes pressure or tension in the energy–momentum tensor, giving it a structure similar to a perfect fluid, i.e., a fluid of strings (FS) Letelier (1983).
-In this connection, recently an equation of state $\rho(r)=\alpha(r)p(r)$ was proposed Nunes dos Santos (2025) for a fluid of strings, leading to a regular black hole model.
-Following the definition of the conserved energy associated with a timelike Killing vector given in Ref.
-Aoki et al.
-(2021), we note that this conserved energy diverges for the standard cloud of strings energy density.
-Motivated by this, we first introduce in this work a modified version of the string-cloud energy density within the fluid-of-strings framework.
-This modification can be interpreted as a geometrical screening of the original distribution.
-As a result, this leads to the convergence of the conserved energy charge.
-In this way, we first provide a new fluid-of-strings black hole model, constructed through the equation of state $\rho(r)=\alpha(r)p(r)$.
-Furthermore, we show that the Ricci scalar near the origin behaves in a way consistent with an integrable singularity, and that the new FS model is also free of a potentially unstable inner horizon.
-On the other hand, the Schwarzschild solution, which according to Birkhoff’s theorem Hawking and Ellis (2023) is the unique spherically symmetric vacuum solution of Einstein’s equations, possesses a strong central singularity.
-As pointed out in reference Ovalle (2024), the Schwarzschild black hole features an event horizon, which can be viewed as a membrane of radius $h$ separating two causally disconnected space-time regions: an interior and an exterior one.
-The interior region, defined by $r<h$, contains the central singularity shielded by the event horizon.
-In this domain, the metric function $f(r)$ becomes negative, causing the radial and temporal coordinates to interchange their roles and thereby revealing the dynamical character of the interior geometry.
-As a consequence, gravitational collapse toward the central singularity becomes unavoidable, in agreement with Penrose’s theorem Penrose (1965).
-Conversely, an observer in the exterior region has no causal access to the interior of the black hole: once the boundary at $r=h$ is crossed, it is impossible to return or communicate with the exterior.
-In this line of research, remarkably, the idea that an interior region, rather than a pointlike mass, can generate a Schwarzschild black hole exterior has recently gained attention Ovalle (2024); Maeda (2025).
-This configuration is obtained by matching an interior solution to the Schwarzschild exterior at the event horizon, without introducing a lightlike thin shell.
-Reference Ovalle (2024) examines whether an alternative structure, beyond a pointlike mass, can give rise to the Schwarzschild black hole region for $r\geq h$.
-In particular, the authors aim to describe an interior configuration associated with this region while keeping the exterior Schwarzschild geometry unchanged and ensuring that tidal forces remain finite everywhere.
-In connection with the latter, the authors propose that the internal region may contain either a central singularity or a regular core.
-Reference Maeda (2025) constructs a so-called fake Schwarzschild model, observationally indistinguishable from the Schwarzschild black hole, by matching a dynamical spherically symmetric interior containing matter to the Schwarzschild exterior at the event horizon without introducing a lightlike thin shell.
-It is shown that this configuration can be obtained from a perfect-fluid solution satisfying a linear equation of state of the type considered in Ref.
-Maeda (2023), analytically continued inside the horizon.
-Reissner Nordström solution represents an electrically charged BH, arising from electromagnetic sources in the energy momentum tensor and characterized by the presence of an inner horizon and a central singularity.
-Motivated by the considerations described in this introduction, and by the existence of an inner horizon and a non-integrable singularity in the Reissner–Nordström solution (RN), in this work, we study the physical and thermodynamic properties of a scenario in which the RN BH possesses an interior region with an integrable singularity and no inner horizon, instead of a pointlike mass.
-We investigate the possibility that the nature of the matter sources capable of supporting such an internal geometry with an integrable singularity corresponds to those associated with clouds of strings and with the new string fluid introduced in this work.
-To carry out this work, we will first define the conditions that a general internal geometry must satisfy in order to represent an integrable singularity, as well as the Israel Darmois Israel (1966) junction conditions between a general internal region and a general external BH geometry.
-We will then examine the connection between the junction conditions and the thermodynamic properties of BHs, and explore their relation to the existence of phase transitions.
-In this context, we will show that satisfying the junction conditions leads to the requirement that the continuity of temperature must hold at the interface, while discontinuities in the tangential pressure signal phase transitions.
-It is worth mentioning that these general conditions could also be of physical interest for future studies of other BH exterior regions, instead of RN, that possess a central singularity, for example BHs with cosmic void density profiles Lustosa et al.
-(2025), BHs surrounded by dark matter Xu et al.
-(2018), hairy BHs Ovalle et al.
-(2021), and quintessential BHs Kiselev (2003); Estrada et al.
+A natural generalization of this model consists of considering a fluid of strings, in which the energy–momentum tensor includes an effective pressure or tension, adopting a structure similar to that of a perfect fluid Letelier (1983).
+Recently, this framework has been further generalized through the introduction of new equations of state for string fluids Nunes dos Santos (2025).
+However, when analyzing the conserved energy associated with timelike Killing vectors, it can be observed that the standard energy density of a cloud of strings leads to divergences in the conserved charge defined in Aoki et al.
 (2021).
-In this way, applying the general conditions described above, in this work we focus on the physical and thermodynamic properties of the mentioned configuration where the exterior geometry is given by the Reissner Nordström BH, while the interior geometry is described by a cloud and by a new fluid of strings with integrable singularities.
-Specifically, we will determine the physical properties that both the cloud of strings (CS) and the new fluid of strings (FS) must satisfy in order to represent the interior of an exterior RN BH region.
-Using the general analysis described above, we will determine the relations between the parameter spaces of both interior regions and the RN exterior BH for which the matching conditions are satisfied and, as a consequence, the temperature remains continuous at the interface.
-In addition, for both interior regions we will determine the conditions relating them to the RN exterior region that must be satisfied for phase transitions to occur or not occur.
-II A new Fluid of strings black hole model with an integrable singularity
+This feature suggests that the effective description of string distributions should be reconsidered when studying physically well-defined gravitational configurations.
+Motivated by this observation, in this work we introduce a modified version of the string-cloud energy density within the framework of a fluid of strings.
+This modification can be interpreted as a geometrical screening of the original distribution and leads to a finite conserved energy.
+On the other hand, it is well known that the Schwarzschild black hole corresponds to a vacuum spacetime endowed with a central singularity.
+Traditionally, this solution is interpreted as being generated by a pointlike mass located at the origin.
+However, recent works have explored the possibility that the exterior Schwarzschild geometry may instead arise from more general internal configurations, obtained by matching an interior spacetime to the Schwarzschild exterior at the event horizon Ovalle (2024); Maeda (2025).
+In this type of construction, the horizon can be interpreted as a surface separating two causally disconnected regions of spacetime, allowing the internal structure to be studied without modifying the geometry observed in the exterior region.
+This perspective naturally raises important physical questions regarding the nature of the matter capable of generating such an interior region.
+In particular, it is natural to ask whether similar configurations could also give rise to non-vacuum black hole geometries, such as electrically charged solutions, and whether an appropriate internal structure could simultaneously address the presence of a central singularity as well as the instabilities and loss of predictability typically associated with spacetimes containing an inner horizon.
+In this context, it is natural to ask whether a similar configuration can be applied to charged solutions, such as the Reissner–Nordström (RN) black hole.
+This solution describes an electrically charged black hole and is characterized by the simultaneous presence of a central singularity and an inner horizon.
+The existence of this inner horizon introduces the well-known problems of instability and loss of predictability discussed above.
+Motivated by these considerations, in this work we investigate a scenario in which the exterior Reissner–Nordström geometry is generated by an extended interior region containing an integrable singularity and lacking an inner horizon.
+To construct this scenario, we first analyze the general conditions that an interior geometry with an integrable singularity must satisfy, as well as the Israel–Darmois matching conditions between an interior region and an exterior black hole solution Israel (1966).
+We further show that these junction conditions admit a natural thermodynamic interpretation: the temperature must remain continuous across the matching surface, while discontinuities in the tangential pressure can be interpreted as signals of phase transitions.
+In order to investigate the nature of the matter capable of generating such an interior region for a Reissner–Nordström black hole, we study configurations in which the interior spacetime is described either by a cloud of strings or by the new fluid-of-strings model introduced in this work, while the exterior region corresponds to the Reissner–Nordström solution.
+In particular, we analyze the physical conditions that lead to the presence of an integrable singularity and the absence of an inner horizon.
+Finally, we determine the relations between the parameters of the interior string distributions and the exterior Reissner–Nordström geometry for which the matching conditions are satisfied and the temperature remains continuous at the interface.
+In addition, we identify the regions of the parameter space for which phase transitions at the event horizon may occur or be absent.
+II A new representation of a fluid of strings black hole model featuring an integrable singularity and finite conserved energy
 In this section, we first present a new string fluid model, showing that it can represent a black hole model over the entire domain $r\in[0,\infty)$.
@@
 III Black Hole models with an Interior Geometry Featuring an Integrable Singularity
+We first analyze the physical configuration of generic matter models for both the interior and exterior regions.
+In the subsequent sections, we focus on the specific case in which the interior region is described by clouds and fluids of strings, while the exterior region is given by the Reissner–Nordström geometry.
+It is worth mentioning that the general framework presented in this section could also be of physical interest for future studies of other black hole exterior geometries, beyond RN, that possess a central singularity.
+Examples include black holes with cosmic void density profiles Lustosa et al.
+(2025), black holes surrounded by dark matter Xu et al.
+(2018), hairy black holes Ovalle et al.
+(2021), and quintessential black holes Kiselev (2003); Estrada et al.
+(2025).
 In Fig.
@@
 (23) and (24).
-IV INTERIOR REGION AS AN INTEGRABLE SINGULARITY
+IV Representation of the INTERIOR REGION AS AN INTEGRABLE SINGULARITY
 In this section, we state the properties that a generic interior geometry must satisfy in order to represent an integrable singularity near the origin.
@@
 It is worth mentioning that this condition leads to the absence of an inner horizon.
-V Junction conditions
+V Junction Conditions and Their Thermodynamic Implications
+In this section, we first analyze the junction conditions between generic matter models describing both the interior and exterior regions.
+In the subsequent sections, we focus on the specific cases of interest in this work, in which the interior region is described by clouds and fluids of strings, while the exterior region is given by the Reissner–Nordström geometry.
+It is worth mentioning that the general junction conditions, together with their physical and thermodynamic implications presented in this section, could also be of interest for future studies of other black hole exterior geometries, beyond RN, that possess a central singularity, as in the examples mentioned in Section III Lustosa et al.
+(2025); Xu et al.
+(2018); Ovalle et al.
+(2021); Kiselev (2003); Estrada et al.
+(2025).
 To construct a model describing a physically viable compact object, it is necessary to impose a smooth matching between the interior geometry and the exterior spacetime.
@@
 On the other hand, using $C=T\,dS/dT=-\,T\,d^{2}G/dT^{2}$, we observe that a discontinuity in the tangential pressure implies the existence of a second-order phase transition.
-VI Example 1: A Reissner–Nordström Exterior with a Cloud of Strings as the Interior Region with an Integrable Singularity
+VI Scenario 1: A Reissner–Nordström Exterior with a Cloud of Strings as the Interior Region with an Integrable Singularity
 In Section  III, we described our geometric setup to represent both the interior and exterior regions of a generic black hole geometry.
@@
 Conversely, for $Q^{2}=Q_{c}^{2}$, no phase transition occurs at this location.
-VII Example 2: A Reissner–Nordström Exterior with a new fluid of Strings as the Interior Region with an Integrable Singularity
+VII Scenario 2: A Reissner–Nordström Exterior with a new fluid of Strings as the Interior Region with an Integrable Singularity
 In Section II, a new string fluid model was introduced.
@@
 VIII Discussion and Summary
-In this work, motivated by the existence of an inner horizon and a non-integrable singularity in the Reissner–Nordström solution (RN), we study the physical and thermodynamic properties of a scenario in which the RN black hole possesses an interior region with an integrable singularity and no inner horizon, instead of a pointlike mass.
-We have investigated the possibility that the nature of the matter sources capable of supporting such an internal geometry with an integrable singularity (IS) corresponds to clouds of strings (CS) and fluids of strings (FS).
-In this connection, motivated by the fact that the conserved charge energy in the CS case diverges, we have proposed a model for the internal region based on an FS in which the energy density, which can be interpreted as a geometrical screening of the original CS distribution, leads to a finite conserved charge energy.
-We have shown that both the CS and FS models used in this work can represent an interior region with an IS and that they lack an inner horizon where predictability would be lost.
-To carry out this work, in Section IV we have determined the conditions that a general internal geometry must satisfy in order to represent an integrable singularity.
-In Section V we have defined the Israel–Darmois Israel (1966) junction conditions between a general internal region and a general external BH geometry.
-In this context, we have shown that satisfying the junction conditions requires the continuity of temperature at the interface, while discontinuities in the tangential pressure signal the presence of phase transitions.
-It is worth mentioning that these general conditions are also of interest for future studies of other BH exterior regions that possess a central singularity instead of the RN one, for example BHs with cosmic void density profiles Lustosa et al.
-(2025), BHs surrounded by dark matter Xu et al.
-(2018), hairy BHs Ovalle et al.
-(2021), and quintessential BHs Kiselev (2003); Estrada et al.
+In this work, we have investigated an alternative scenario for the internal structure of the Reissner–Nordström (RN) black hole, motivated by the simultaneous presence of a central non-integrable singularity and an inner horizon acting as a Cauchy horizon, which is commonly associated with instabilities and a potential loss of predictability in General Relativity.
+We have shown that the exterior RN geometry can instead be generated by an extended interior region containing an integrable singularity and lacking an inner horizon, thus replacing the conventional description based on a pointlike mass located at the origin.
+We have analyzed the general conditions required to match an interior region with an exterior black hole geometry through the Israel–Darmois junction conditions Israel (1966).
+We show that thermodynamic consistency across the interface requires the continuity of temperature, suggesting that the two regions must be in thermal equilibrium, while discontinuities in the tangential pressure can signal the presence of gravitational phase transitions.
+This thermodynamic interpretation provides an additional physical perspective on the matching between the interior and exterior geometries.
+To explore the physical nature of the matter capable of sustaining such an interior geometry while generating a Reissner–Nordström exterior geometry, we considered two explicit realizations based on string distributions: a cloud of strings (CS) and a fluid of strings (FS).
+In the case of the cloud of strings, following the definition of conserved energy given in Ref.
+Aoki et al.
+(2021), the analysis of the conserved quantities associated with timelike Killing vectors reveals that the conserved energy diverges.
+Motivated by this feature, we introduce a new fluid-of-strings (FS) model in which the energy density can be interpreted as a geometrically screened version of the standard string-cloud distribution.
+This modification leads to configurations that, when extended to infinity, represent black holes with finite conserved energy, in contrast to the divergent conserved energy typically encountered in the string-cloud black hole.
+In this way, for both interior regions constructed from the cloud-of-strings and the new fluid-of-strings distributions, we analyze the physical conditions that lead to the presence of an integrable singularity and the absence of an inner horizon.
+Furthermore, we determine the relations between the parameters of the two interior string distributions and the exterior Reissner–Nordström geometry for which the temperature remains continuous at the interface, thereby implying thermal equilibrium across the matching surface.
+In addition, we identify the regions of parameter space for which phase transitions at the event horizon may occur or be absent.
+It is worth mentioning that the framework developed in this work is not restricted to the Reissner–Nordström geometry.
+The general conditions obtained for interior geometries containing integrable singularities, together with their thermodynamic implications, may also be relevant for future studies of other black hole exterior spacetimes that exhibit a destructive central singularity and, in some cases, an inner horizon.
+In such situations, these exterior geometries could potentially be generated by an interior region endowed with an integrable singularity and lacking an inner horizon.
+Examples of black hole configurations that could be investigated within the framework presented here include black holes with cosmic void density profiles Lustosa et al.
+(2025), black holes surrounded by dark matter Xu et al.
+(2018), hairy black holes Ovalle et al.
+(2021), and quintessential black holes Kiselev (2003); Estrada et al.
 (2025).
-Using the framework developed in the previous paragraph, in Sections VI and VII we studied the cases in which the internal geometry of the Reissner–Nordström spacetime is described, respectively, by a string cloud and by our newly proposed string fluid model.
-In both scenarios, the parameters of the corresponding internal geometries satisfy the conditions listed in Section IV, ensuring the presence of an integrable singularity near the origin, the absence of a potentially unstable de Sitter core, and the absence of an internal horizon that could lead to a breakdown of causality.
-In Section VI, we determined the allowed values of the energy density parameter $a$ and of the internal cosmological constant required to satisfy both Israel junction conditions at the interface between the external RN region and the internal string cloud region.
-In addition, we identified a critical value of the electric charge for which no phase transition occurs in the location of the interface between the two regions.
-In Section VII, we derived the conditions that the AdS radius, the electric charge, the screening parameter $b$, and the event horizon must satisfy in order to fulfill both Israel junction conditions at the interface between the external RN region and the internal region described by the new string fluid.
-Moreover, we numerically determined the critical value of the screening parameter for which no phase transition occurs in the location of the interface between the two regions.
 Appendix A A brief revision of the cloud of string model